### v1

# Generative Anchored Fields: Controlled Data Generation via Emergent Velocity Fields and Transport Algebra

Abstract
We present Generative Anchored Fields (GAF), a generative model that learns independent endpoint predictors $J$ (noise) and $K$ (data) rather than a trajectory predictor. The velocity field v=KJv=K-J emerges from their time-conditioned disagreement. This factorization enables Transport Algebra: algebraic operation on learned $\{(J_{n},K_{n})\}_{n=1}^{N}$ heads for compositional control. With class-specific $K_{n}$ heads, GAF supports a rich family of directed transport maps between a shared base distribution and multiple modalities, enabling controllable interpolation, hybrid generation, and semantic morphing through vector arithmetic. We achieve strong sample quality (FID 7.5 on CelebA-HQ $64\times 64$) while uniquely providing compositional generation as an architectural primitive. We further demonstrate, GAF has lossless cyclic transport between its initial and final state with LPIPS=$0.0$. Code available at https://github.com/IDLabMedia/GAF
1 Introduction
Modern generative models [kingma_auto-encoding_2013, goodfellow_generative_2014, ho_denoising_2020] achieve remarkable sample quality but lack precise compositional control. Score-based Diffusion and flow matching models learn a single time-conditioned predictor that maps noise to data via a unified field [pmlr-v37-sohl-dickstein15, lipman_flow_2023]. Although models such as Stable Diffusion [Rombach_2022_CVPR, podell2024sdxl] and FLUX.1 [blackforestlabs2025flux] generate images with high fidelity, their monolithic architecture treats control as an external steering process, implemented through classifier-free guidance [ho2021classifierfree], prompt engineering, latent space editing [mokady2022null], or attention manipulation [hertz2022prompt], rather than an intrinsic property of the learned representation.
Figure 1: Generated samples demonstrating our transport algebra framework. (a) Selected samples from CelebA-HQ ($256\times 256$ px), demonstrating high-fidelity generation across diverse demographics and attributes. (b) Lossless cyclic transport across three domains: $v_{\text{cat}_{0}}\to v_{\text{dog}}\to v_{\text{wild}}\to v_{\text{cat}_{0}}$ with LPIPS=$0.0$. (c) Barycentric interpolation: $v_{i\to j\to k}=\alpha v_{i}+\beta v_{j}+\gamma v_{k}$, showing smooth multi-domain transitions in velocity space.
These limitations suggest a need to re-examine the core learning objectives. The trajectory-based paradigm focuses on learning step-by-step dynamics along a path [liu2022flow]. This raises a fundamental question: must we learn a step-by-step mechanism of the trajectory or is it sufficient to know only its origin and destination? The trajectory-based paradigm assumes the first option. We explore the second one.
We present Generative Anchored Fields (GAF), where generation is an explicit algebraic operation. Rather than learning a single trajectory predictor, GAF learns independent endpoint operators $\{(J_{n},K_{n})\}_{n=1}^{N}$ that can be linearly combined to achieve precise compositional control of the generative process. Built on the principle of endpoint knowledge, the model uses a shared trunk, $\Phi$, which produces a time-conditioned feature bank to feed a pair of twin predictors, $J$ and $K$. The twins have opposing objectives: one is anchored to the noise distribution, the other to the data manifold. Unlike trajectory-based methods, GAF does not learn the velocity directly. Instead, the velocity field is an emergent consequence of the twins’ time-conditioned disagreement. Put simply, GAF is not trained to follow a path; it is trained only to know its endpoints.
This factorization has a direct consequence: precise and controllable generation. Because GAF learns independent endpoint predictors $\{(J_{n},K_{n})\}_{n=1}^{N}$, particularly when using class-specific $K_{n}$ heads, it naturally enables compositional operations at inference time. One can switch between different $K$ heads mid-trajectory, interpolate between class manifolds with explicit geometric control, or arithmetically combine predictors to generate novel outputs. These operations are architectural primitives in GAF. We term this capability Transport Algebra: the ability to perform algebraic manipulations directly on the model’s learned components to achieve precise semantic control.
Our proposed model has the following contributions:
  • Transport Algebra. GAF introduces transport algebra to generative modeling: the ability to perform arithmetic operations directly on learned endpoint predictors. This enables controllable test-time generation, such as interpolation between classes, switching manifolds mid-trajectory, or combining multiple classes for novel generation.
  • Emergent dynamics via endpoint-anchored learning. Instead of directly learning a score, noise or velocity field, GAF trains two time-conditioned endpoint predictors $J$ and $K$ that target a noise and data endpoint, respectively. The transport velocity field $v(\mathbf{x}_{t},t)=K-J$ emerges as their antisymmetric disagreement. This decouples the training objective (endpoint accuracy) from the sampler, which performs Ordinary Differential Equation (ODE) integration.
[…]
  • Inherent modularity and extensibility. GAF’s shared trunk with pluggable “twin” or “tuplet” heads is inherently modular. One can easily add additional $K_{n}$ heads without altering the trunk.
  • Simple training and deterministic sampling. Training is a direct endpoint regression with simple residual and swap regularizers. Sampling is deterministic and guidance free, requiring only single-step Euler integration due to GAF’s self-correcting dynamics, unlike multi-step ODE solvers needed by other flow-based models..
2 Related Work
Modern DGM is largely dominated by methods that learn to reverse a data-to-noise trajectory, with key differences arising in the parametrization of this reverse process [song2021scorebased, albergo2023building, lipman_flow_2023]. These approaches can be broadly categorized into score-based, denoising and velocity-based formulations. Our work builds upon the velocity-based formalism of generative processes but introduces a novel endpoint regression objective for the underlying transport dynamics.
Score-based generative models learn the score function (i.e., the gradient of the log-density $\nabla_{\mathbf{x}}\log p(\mathbf{x}_{t})$) of noise-perturbed data distribution [10.5555/3454287.3455354, 10.5555/3495724.3496767]. By training a noise conditional score network across multiple noise scales, these models can generate samples via annealed Langevin dynamics. This paradigm was later formalized through the lens of stochastic differential equations (SDE), which established the score function as the key component needed to reverse a continuous-time forward diffusion process [song2021scorebased]. This SDE framework provided a unifying mathematical language for a wide variety of diffusion models, but the core objective remained the estimation of the score.
Concurrently, Denoising Diffusion Probabilistic Models (DDPMs) [pmlr-v37-sohl-dickstein15, ho_denoising_2020] achieved state-of-the-art sample quality by simplifying the transport objective to predict the noise $\epsilon$ added at each step of a discrete forward process. This noise-prediction ($\epsilon$-prediction) target proved highly stable and effective, becoming the default standard, such as in OpenAI’s DALL-E [dhariwal_diffusion_2021], Google’s Imagen [saharia_photorealistic_2022], and Midjourney [pmlr-v162-nichol22a]. Later, Denoising Diffusion Implicit Models (DDIM) [song2021denoising] introduced deterministic sampling for diffusion models to remove randomness during generation, allowing the model to skip steps and produce high-quality images. Subsequent improvements focused on optimization, the training recipe, and architecture, with new methods like cosine noise schedules, learned variances [nichol_improved_2021], architectural upgrades [dhariwal_diffusion_2021], and powerful conditioning mechanisms [ho2021classifierfree]. A pivotal development was classifier-free guidance (CFG) [ho2021classifierfree], which enabled high-fidelity conditional generation from a single model by extrapolating between conditional and unconditional outputs. While our work uses the powerful transformer backbone (DiT) [vaswani_attention_2017, dosovitskiy2021an, williampeebles2023] that pioneered in this domain, we depart from the standard practice of predicting either the score or the noise.
Most recently, the field has increasingly shifted towards a deterministic, continuous-time perspective rooted in ODEs and optimal transport [villani2009optimal, lipman_flow_2023, liu2022flow, lipman2024flowmatchingguidecode]. These models learn a velocity field $v_{\theta}(\mathbf{x}_{t},t)$ that governs a probability flow ODE, transporting samples from a simple base distribution (such as a Gaussian distribution) to the data distribution. While mathematically related to the score, this velocity ($v$-prediction) parametrization was shown to have better numerical stability and scaling properties [salimans2022progressive, 10.5555/3600270.3602196]. The development of simulation-free training objectives, such as Flow Matching [lipman_flow_2023] and Rectified Flows [liu2022flow], has made this approach highly efficient. By directly regressing the network’s output onto a target field, often defining a straight path between noise and data, these methods learn linear trajectories, enabling high-quality generation in very few sampling steps, as demonstrated by models like FLUX.1 [blackforestlabs2025flux], Stable Diffusion 3 [esser_scaling_2024], and InstaFlow [ICLR2024_4dc37a7b].
All these approaches, whether predicting score, noise, or velocity, learn a single time-conditioned predictor. While effective for generation, this monolithic architecture limits compositional control: one cannot independently manipulate learned representations for different classes or combine them arithmetically. Conditional generation instead relies on guidance mechanisms like CFG or prompt engineering.
Recent work has explored alternative control mechanism. For example, Compositional Visual Generation [liu2022compositional] defines two compositional operators (AND and NOT) for multi-concept generation, using a set of diffusion models to enabling logical operation over concepts. Moreover, ControlNet [10377881] enables spatial control in text-to-image generation by locking the original diffusion model while training another network for conditional control with extra inputs like edges or depth. Although these methods provide impressive control, they require either maintaining multiple models for composing their outputs or training additional conditioning networks. Both approaches treat control as an add-on mechanism rather than an intrinsic architectural property.
To address this limitation, GAF introduces an algebraic factorization of the transport dynamics. Instead of learning a single, monolithic velocity field $v_{\theta}$, GAF learns two endpoint predictors: a twin $J$ that regresses towards the noise distribution and a twin $K$ that regresses towards the data manifold. The velocity field $v$ then emerges from their difference, v=KJv=K-J, enforced by a paired and time-antisymmetric loss that promotes consistency between the forward and backward (reverse in time) dynamics. This explicit factorization of the transport operators ($J$ and $K$) enables unique architectural choices, such as independent, class-specialized $K$-heads. Unlike external composition [liu2022compositional] or auxiliary networks [10377881], GAF’s compositional capabilities are intrinsic: they emerge from the factorized architecture. This design unlocks novel inference-time capabilities, allowing vector operations between class distributions, a process we now term Transport Algebra. This form of algebraic manipulation over $\{(J_{n},K_{n})\}_{n=1}^{N}$ to generate new samples with precise compositional control is not naturally afforded by standard score-, flow-, or noise-based conditional generative models.
3 Generative Anchored Fields (GAF)
3.1 The GAF Framework
We introduce the Generative Anchored Fields model. Let $\mathbf{z}_{x},\mathbf{z}_{y}\in\mathbb{R}^{d}$ be samples from a data distribution $\mathbf{z}_{x}\sim p_{\text{data}}(\mathbf{x})$ and a standard normal distribution $\mathbf{z}_{y}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, respectively. Additionally, let $\mathbf{c}\in{\mathbb{R}^{d}}$ be an optional conditioning vector (e.g., a class embedding).
The model operates on linear bridges connecting these samples:
$\mathbf{x}_{t}=(1-t)\mathbf{z}_{y}+t\mathbf{z}_{x},\quad t\in[0,1].$ (1)
The central task is for GAF to learn the two anchor endpoints $\mathbf{z}_{y}$ and $\mathbf{z}_{x}$ from a given point on the bridge that they define. We call this task Coordinate Learning: the ability to identify the coordinates (endpoints) of a bridge from any point along its trajectory. The core of the model is a neural network, referred to as the trunk $\Phi$, that processes the bridge $\mathbf{x}_{t}$ and timestep $t$ to produce a feature bank $\mathbf{f}_{t}$:
$\mathbf{f}_{t}=\Phi(\mathbf{x}_{t},t).$ (2)
Optionally, for class guidance we condition on $\mathbf{c}$:
$\mathbf{f}_{t}=\Phi(\mathbf{x}_{t},t,\mathbf{c}).$ (3)
The feature bank $\mathbf{f}_{t}$ is then processed by a pair of twins $J$ and $K$, which regress the bridge $\mathbf{x}_{t}$ towards its endpoints (i.e., noise $\mathbf{z}_{y}$ and data $\mathbf{z}_{x}$, respectively). The twins are formally defined as:
$J$ $:=(1-t)\mathbf{x}_{t}+H_{J}(\mathbf{f}_{t}),$ (4)
$K$ $:=t\mathbf{x}_{t}+H_{K}(\mathbf{f}_{t}).$ (5)
$H_{J}$ and $H_{K}$ are identical neural network architectures, but trained for separate tasks. We denote $J_{\text{res}}:=H_{J}(\mathbf{f}_{t})$ and $K_{\text{res}}:=H_{K}(\mathbf{f}_{t})$. We term our model “Anchored Fields” because the predictors $J$ and $K$ are anchored to fixed endpoints: $J$ targets the noise distribution, $K$ targets the data manifold. Unlike trajectory-based models that learn the dynamics between these points, GAF anchors its learning directly to the endpoints themselves.
3.1.1 Velocity Field and ODE
The twins move in opposite directions along the bridge. More specifically, $J$ is oriented toward the noise endpoint, and $K$ toward the data endpoint. Their ideal boundary conditions are $J=\mathbf{z}_{y}$ at $t=0$, and $K=\mathbf{z}_{x}$ at $t=1$. To quantify the twins’ motion relative to each other along the bridge, we compute the path’s instantaneous velocity with respect to $t$. For the linear bridge (1), we have:
$\frac{d\mathbf{x}_{t}}{dt}=\frac{d}{dt}\big[(1-t)\mathbf{z}_{y}+t\mathbf{z}_{x}\big]=\mathbf{z}_{x}-\mathbf{z}_{y}.$ (6)
With this ground-truth velocity in hand, we define the learned velocity field $v$ as the twins disagreement:
$v(\mathbf{x}_{t},t)=K-J.$ (7)
This learned field then specifies an ODE that transports the noise sample $\mathbf{z}_{y}$ to the data sample $\mathbf{z}_{x}$:
$\frac{d\mathbf{{x}}_{t}}{dt}=v(\mathbf{x}_{t},t).$ (8)
We term $J$ and $K$ twins to emphasis their symmetric roles: both are endpoint predictors that must be learned in balance. Their velocity field v=KJv=K-J emerges from their difference, so imbalanced training would bias their dynamics towards one endpoint.
We generate new data samples by integrating this ODE from $t=0$ to $t=1$ along either a linear or cosine time grid, using a single-step Euler integration [chen_neural_2018, 10.5555/3600270.3602196].
3.1.2 GAF Training
The model is trained using the twins pair loss:
$\mathcal{L_{\text{pair}}}=\mathbb{E}_{\mathbf{z}_{x},\mathbf{z}_{y},t}[(1-t)\|J-\mathbf{z}_{y}\|_{2}^{2}+t\|K-\mathbf{z}_{x}\|_{2}^{2}]$ (9)
To improve training stability, we also enforce the model’s boundary behavior. That is, we examine the prediction errors at the ideal boundary condition $J-\mathbf{z}_{y}$ and $K-\mathbf{z}_{x}$ by expanding them at the endpoints.
From $J=(1-t)\mathbf{x}_{t}+J_{\text{res}}$, at $t=0$, we have $J=\mathbf{z}_{y}+J_{\text{res}}$. Thus, achieving the ideal condition $J=\mathbf{z}_{y}$ requires $J_{\text{res}}=0$. By substituting the definition of $J$, we get:
$J=\mathbf{z}_{y}\equiv 0$ $=J-\mathbf{z}_{y}$ (10)
$=(1-t)\mathbf{x}_{t}+J_{\text{res}}-\mathbf{z}_{y}$
$=(1-t)((1-t)\mathbf{z}_{y}+t\mathbf{z}_{x})+J_{\text{res}}-\mathbf{z}_{y}$
$=((1-t)\mathbf{z}_{y}+t\mathbf{z}_{x})-(t(1-t)\mathbf{z}_{y}-t^{2}\mathbf{z}_{x})+J_{\text{res}}-\mathbf{z}_{y}$
$=(1-t)\mathbf{z}_{y}+t\mathbf{z}_{x}-t^{2}\mathbf{z}_{x}-t(1-t)\mathbf{z}_{y}+J_{\text{res}}-\mathbf{z}_{y}$
$\equiv\underbrace{t(1-t)(\mathbf{z}_{x}-\mathbf{z}_{y})}_{\text{cross term}}-\underbrace{t\mathbf{z}_{y}}_{\text{endpoint}}+\underbrace{J_{\text{res}}}_{\text{residual}}$ (11)
Since our ideal learning goal is to have $J=\mathbf{z}_{y}$ at $t=0$, the ideal boundary condition implies the magnitude of the learned residual term $J_{\text{res}}$ must be zero at this point: $J_{\text{res}}|_{t=0}=0$.
Similarly, from $K=t\mathbf{x}_{t}+K_{\text{res}}$, at $t=1$, $K=\mathbf{z}_{x}+K_{\text{res}}$. Thus, achieving the ideal condition $K=\mathbf{z}_{x}$ requires $K_{\text{res}}=0$. By substituting the definition of $K$, we get:
$K=\mathbf{z}_{x}\equiv 0$ $=K-\mathbf{z}_{x}$ (12)
$=t\mathbf{x}_{t}+K_{\text{res}}-\mathbf{z}_{x}$
$=t((1-t)\mathbf{z}_{y}+t\mathbf{z}_{x})+K_{\text{res}}-\mathbf{z}_{x}$
$=t(1-t)\mathbf{z}_{y}+t^{2}\mathbf{z}_{x}-\mathbf{z}_{x}+K_{\text{res}}$
$=t(1-t)\mathbf{z}_{y}+\mathbf{z}_{x}(t^{2}-1)+K_{\text{res}},\quad t^{2}-1=(1-t)(-1-t)$
$=t(1-t)\mathbf{z}_{y}+\mathbf{z}_{x}(1-t)(-1-t)+K_{\text{res}}$
$=(1-t)(t\mathbf{z}_{y}+\mathbf{z}_{x}(-1-t))+K_{\text{res}}$
$=(1-t)(t\mathbf{z}_{y}-\mathbf{z}_{x}-t\mathbf{z}_{x})+K_{\text{res}}$
$=(1-t)(t(\mathbf{z}_{y}-\mathbf{z}_{x})-\mathbf{z}_{x})+K_{\text{res}}$
$=t(1-t)(\mathbf{z}_{y}-\mathbf{z}_{x})-(1-t)\mathbf{z}_{x}+K_{\text{res}}$
$\equiv\underbrace{t(1-t)(\mathbf{z}_{y}-\mathbf{z}_{x})}_{\text{cross term}}-\underbrace{(1-t)\mathbf{z}_{x}}_{\text{endpoint}}+\underbrace{K_{\text{res}}}_{\text{residual}}$ (13)
Since our ideal learning goal is to have $K=\mathbf{z}_{x}$ at $t=1$, the ideal boundary condition implies the magnitude of the learned residual term $K_{\text{res}}$ must be zero at this point: $K_{\text{res}}|_{t=1}=0$.
In both derivations in (11) and (13), the cross term represents the interaction between the endpoints. The factor t(1t)t(1-t) acts as a symmetric weighting or modulation factor. Its weight peaks at the midpoint, $t=\frac{1}{2}$, concentrating the error signal at this point of maximum complexity. This forces both $J$ and $K$ to prioritize learning the correct transition path, rather than just matching the endpoints themselves.
While the primary endpoint loss (9) implicitly encourages the ideal boundary conditions, we introduce an explicit residual regularization penalty to enforce them directly and improve stability. From (11) and (13), we have our residual penalty as:
$\mathcal{L_{\text{res}}}=(1-t)\|J_{\text{res}}\|_{2}^{2}+t\|K_{\text{res}}\|_{2}^{2}$ (14)
The endpoint residual penalty ensures $J\rightarrow\mathbf{z}_{y}$ as $t\rightarrow 0$ and $K\rightarrow\mathbf{z}_{x}$ as $t\rightarrow 1$, but it does not fully determine the intermediate behavior of GAF on the bridge. Since we want GAF to behave consistently across the entire bridge, we need a mechanism to do that. One way to do that is by swapping the residuals $J_{\text{res}}$ and $K_{\text{res}}$ across the bridge and create antisymmetric relationship between them. We define the swap operator (function) acting on the residuals at $t$ and 1t1-t as:
$\mathcal{S}:(\mathbf{z}_{x},\mathbf{z}_{y},t,c)\mapsto(\mathbf{z}_{y},\mathbf{z}_{x},1-t,c)$ (15)
Using our bridge formalism in (1) and the swap operator in (15), we have:
$\mathbf{x}_{1-t}$ $=(1-(1-t))\mathbf{z}_{x}+(1-t)\mathbf{z}_{y}$ (16)
$\mathbf{x}_{1-t}$ $=t\mathbf{z}_{x}+(1-t)\mathbf{z}_{y}$ (17)
From (1) and (17), $\mathbf{x}_{t}\stackrel{{\scriptstyle\mathcal{S}}}{{=}}\mathbf{x}_{1-t}$, i.e., the swap stays on the bridge, and only the direction of travel changes.
Here, $\stackrel{{\scriptstyle\mathcal{S}}}{{=}}$ means equality after applying the swap (namely, swapping endpoints and flipping time).
Using our trunk from (3) and swap function from (15), the endpoint-swap antisymmetric residuals are defined as:
$\tilde{J}_{\text{res}}:=H_{J}(\Phi(\mathbf{x}_{1-t},1-t,c)),\quad\tilde{K}_{\text{res}}:=H_{K}(\Phi(\mathbf{x}_{1-t},1-t,c)).$ (18)
Similarly, we have:
$v(t)$ $=t\mathbf{x}_{t}+K_{\text{res}}-((1-t)\mathbf{x}_{t}+J_{\text{res}})$ $v(1-t)$ $=(1-t)\mathbf{x}_{1-t}+\tilde{K}_{\text{res}}-((1-(1-t))\mathbf{x}_{1-t}+\tilde{J}_{\text{res}})$
$=(2t-1)\mathbf{x}_{t}+(K_{\text{res}}-J_{\text{res}})$ $=(1-2t)\mathbf{x}_{1-t}+(\tilde{K}_{\text{res}}-\tilde{J}_{\text{res}})$ (19)
Our residual antisymmetric targets are:
$J_{\text{res}}=-\tilde{K}_{\text{res}},\quad K_{\text{res}}=-\tilde{J}_{\text{res}}\;\Rightarrow\;v(1-t)=-v(t)$ (20)
Thus, the swap function flips the velocity field in time.
Although GAF is trained with an endpoint regression, its emergent velocity field exhibits how it learns the transport dynamics. From (19), we have $v(\mathbf{x}_{t},t)=(2t-1)\mathbf{x}_{t}+(K_{\text{res}}-J_{\text{res}})$. Let $\Delta_{\text{res}}=K_{\text{res}}-J_{\text{res}}$ and examine $v$ at $t\in[0,1]$.
$v(\mathbf{x}_{t},t)=\begin{cases}-\mathbf{z}_{y}+\Delta_{\text{res}}&\text{lim}_{t\rightarrow 0}\\ \Delta_{\text{res}}&t=\frac{1}{2}\\ \mathbf{z}_{x}+\Delta_{\text{res}}&\text{lim}_{t\rightarrow 1}\end{cases}$ (21)
Similarly, from the swap operator in (19), we have $v(\mathbf{x}_{1-t},t)=(1-2t)\mathbf{x}_{1-t}+(\tilde{K}_{\text{res}}-\tilde{J}_{\text{res}})$. Let $\tilde{\Delta}_{\text{res}}=\tilde{K}_{\text{res}}-\tilde{J}_{\text{res}}$ and examine $v$ at $t\in[0,1]$.
$v(\mathbf{x}_{1-t},t)=\begin{cases}\mathbf{z}_{y}+\tilde{\Delta}_{\text{res}}&\text{lim}_{t\rightarrow 0}\\ \tilde{\Delta}_{\text{res}}&t=\frac{1}{2}\\ -\mathbf{z}_{x}+\tilde{\Delta}_{\text{res}}&\text{lim}_{t\rightarrow 1}\end{cases}$ (22)
As we have discussed earlier, in GAF, we do not train the velocity field directly, as in rectified flow. Instead, the velocity simply emerges from the twins opposite movement. However, from (21) and (22), we can see that the emergent velocity field incorporates knowledge of the trajectory’s endpoint.
Let $g_{0}=J_{\text{res}}+\tilde{K}_{\text{res}}$ and $g_{1}=K_{\text{res}}+\tilde{J}_{\text{res}}$. The ideal time-antisymmetric condition is $g_{0}=g_{1}=0$. Hence, we define the residual-antisymmetry penalty as:
$\mathcal{L}_{\text{swap}}$ $=\|g_{0}\|_{2}^{2}+\|g_{1}\|_{2}^{2}$
$=\|J_{\text{res}}+\tilde{K}_{\text{res}}\|_{2}^{2}+\|K_{\text{res}}+\tilde{J}_{\text{res}}\|_{2}^{2}.$ (23)
From (9), (14), and (23), the GAF training loss is
$\mathcal{L}_{\text{GAF}}=\mathcal{L}_{\text{pair}}+\lambda_{\text{res}}\,\mathcal{L}_{\text{res}}+\lambda_{\text{swap}}\,\mathcal{L}_{\text{swap}}.$ (24)
where $\lambda_{\text{res}}$ and $\lambda_{\text{swap}}$ are hyperparameters (e.g, both $0.01$).
The overall architecture of GAF and its data flow is summarized in Figure 2.
Figure 2: The high-level overview of GAF architecture and its data flow.
3.2 Properties of GAF
3.2.1 Main Components
GAF comprises four important components. Each component is necessary for the model to retain its defining behavior; removing one of them either destroys the emergent velocity, or reduces GAF to a different formulation or eliminates learning altogether.
  1. 1.
    Trunk (shared backbone). The trunk $\Phi(\mathbf{x}_{t},t)\in\mathbb{R}^{d}$ provides the shared time-conditioned representation used by the twins. It hides architectural complexity details from the twins and enables modularity (plugging different heads, conditioning, or label embeddings) and parameter sharing. Removing the trunk would destroy this modularization ability and prevent the twins from coordinating through a common state. Hence, if you remove the trunk, you would lose the modularity of GAF.
  2. 2.
    Twins (opposing heads). The two time-conditioned predictors $J$, and $K$ pull the anchored state $\mathbf{x}_{t}$ toward noise and data. Their antisymmetric disagreement $v(\mathbf{x}_{t},t)=K-J$ yields the emergent velocity field. Removing either head eliminates the disagreement, the emergent velocity, and, by design, GAF itself.
  3. 3.
    Twin Pair Loss (endpoint anchoring). Each twin is trained with its endpoint reconstruction. The pairs are required to anchor both ends of the path. Dropping one of the losses collapses GAF.
  4. 4.
    Sampler (single ODE solver). Sampling integrates the emergent field with a single ODE solver (e.g., Heun sampler). No auxiliary corrector network, guidance model, or second model is required. Adding an additional learned sampler is possible but would depart from minimal GAF design.
3.2.2 GAF is Naturally Modular
GAF is a naturally modular model. This means that one can simply add additional $J$ and $K$ heads on to the trunk to add new members to the twins or “tuplets”, without changing any other part of the architecture. Each head represents a different noise region and modality. For example, one could simultaneously learn an image with its albedo ($K_{1}$), depth map ($K_{2}$), and normal map ($K_{3}$), with each modality learned by specific member of the twins. We can add tuplets ($K_{n}$) as long as the base trunk is capable of feeding enough features for all tuplets, and as long as it is designed for the specific modality.
GAF is modal-agnostic. That is, to change the modality of GAF, we only need to change the trunk’s modality and the I/O. For example, to change the modality from image data to video data, you only need to change the trunk. The other components need no (or minimal) change.
3.3 Transport Algebra
Transport Algebra organizes GAF’s endpoint pairs into an algebra of emergent velocity fields, enabling composable operations that yield controlled data generation. To understand this, recall that GAF defines two time-conditioned endpoints $J$ and $K$. The velocity field that arises from their disagreement $v(\mathbf{x}_{t},t)=K-J$ defines a time-conditioned vector field. Because GAF is modular, the shared trunk can host multiple tuplets $\{(J_{n},K_{n})\}_{n=1}^{N}$, yielding $v_{n}=K_{n}-J_{n}$ emergent velocity fields. This creates a natural algebra of transport in which vector field operations on $J_{n},K_{n},v_{n}$ correspond to vector combinations of directions in the field. This ability of GAF to operate algebraically leads to a precise and semantically coherent method of data generation across multiple classes and domains.
Bridge and swap operators.
The swap operator from (15) carves a two-way lane on a bridge and enforces a bidirectional highway for transport on the same bridge. Using the swap operator, we define three elementary operators on the bridge configurations for multiple path traversals as follows:
$\mathcal{S}_{\text{swap}}$ $:(\mathbf{z}_{y},\mathbf{z}_{x},t,c)\mapsto(\mathbf{z}_{x},\mathbf{z}_{y},t,c)$ (Swap endpoints) (25)
$\mathcal{S}_{\text{flip}}$ $:(\mathbf{z}_{y},\mathbf{z}_{x},t,c)\mapsto(\mathbf{z}_{y},\mathbf{z}_{x},1-t,c)$ (Flip time) (26)
$\mathcal{S}_{\text{swap \& flip}}$ $:(\mathbf{z}_{y},\mathbf{z}_{x},t,c)\mapsto(\mathbf{z}_{x},\mathbf{z}_{y},1-t,c)$ (Swap & flip) (27)
From (25) and (26), we get $\mathbf{x}_{t}=(1-t)\mathbf{z}_{x}+t\mathbf{z}_{y}$. Just like the main bridge in (15), the new swaps in (25) and (26) likewise remain on a bridge. Consequently, one can compose these operators to perform controlled generation at test time, using the resulting configurations to traverse or switch modalities, see Section 5.5. Figure 3 illustrates the four swap configurations.
Figure 3: Visualization of the swap operator $\mathcal{S}$ acting on a bridge. (A) The initial configuration of the bridge (1). Subfigures (B), (C), and (D) illustrate the results of applying the operations described in equations (25), (26), and (27), respectively.
$J/K$ pairing topologies.
The swap operator $\mathcal{S}$ gives us the mechanism to traverse through the endpoints back and forth in time. In addition, we can also consider the three main different ways multiple pairs of twin $J$ and $K$ can be mapped into a $\{(J_{n},K_{n})\}_{n=1}^{N}$ mappings. The pairings are: (A) one-to-one, (B) star (one-to-many), and (C) clustered mapping, shown in Figure 4.
Figure 4: Three $J/K$ pairing topologies: (A) one-to-one, (B) star (one-to-many), and (C) clustered. Each topology shows different ways to map the noise endpoint $J$ to the data endpoint $K$.
Transport algebra operations.
GAF’s independent $K$ heads enable compositional transport through linear velocity operations. Let $i,j\in\{1,\dots,N\}$.
  1. 1.
    Single-class transport. If the noise endpoint $J$ is shared, then
    $v_{i}=K_{i}-J$
    If each modality has its own noise endpoint $J_{i}$, then:
    $v_{i}=K_{i}-J_{i}$
    Integrating $v_{i}$ from noise ($t=0$) to data ($t=1$) generates samples from class $i$.
  2. 2.
    Cross-modality transport (via $J$). Move between data endpoints of two (or more) domains. To move from class $i$ to class $j$, we use $J$ as a terminal and $v_{i\to j}=K_{i}-K_{j}$ expands to an encode-decode process:
    1. (a)
      Decode to class $i$: Integrate forward with $v_{J\to i}=K_{i}-J$ to reach $K_{i}$ (forward time)
    2. (b)
      Encode to noise: Integrate with $v_{i\to J}=J-K_{i}$ to reach $J$ (reverse time)
    3. (c)
      Decode to target class $j$: Integrate forward with $v_{J\to j}=K_{j}-J$ to reach $K_{j}$ (forward time)
      The full transport path is:
      $v_{i\to j}=K_{i}\xrightarrow{-(K_{i}-J)}J\xrightarrow{K_{j}-J}K_{j}$
    This sequential transport process uses $J$ as a universal intermediate representation shared across all classes, enabling transport between any pair of $K$ heads without requiring direct $K_{i}\leftrightarrow K_{j}$.
  3. 3.
    Direct $K$-head interpolation. We can interpolate directly in the velocity space via weighted combinations
    $v_{\alpha}(\mathbf{x}_{t},t)=(1-\alpha)(K_{i}-J)+\alpha(K_{j}-J)=(1-\alpha)v_{i}+\alpha v_{j},$ (28)
    for $\alpha\in[0,1]$.
    A simultaneous weighted combination of multiple $K$ heads enables compositional generation, which is experimented on in Section 5.5.3. These operations demonstrate GAF’s transport algebra where $K$ independent heads act as linear operators that can be composed, interpolated, and combined to traverse the generative space. This property extends naturally to sequential domains. For example, for video generation, we hypothesize that GAF can smoothly transition between temporal states while preserving semantic coherence through velocity interpolation, which we will explore in future work.
  4. 4.
    Multi-class composition (barycentric). Combine multiple velocity fields by linear vector operations.
    $v_{\text{blend}}=\sum_{m}w_{m}(K_{m}-J),\text{with}\sum_{m}w_{m}=1$
By using $J_{n},K_{n},v_{n}$ that live in $\mathcal{S}$ and the $\{(J_{n},K_{n})\}_{n=1}^{N}$ pairings, one can traverse modalities, switch modalities mid generation, revert to the previous modality, chain back to the origin, or even use a specific generated modality as a bootstrap for a chained generation. All of this is possible by switching between the GAF’s velocity field $v_{i}$, $i\in\{1,\dots,N\}$, and applying simple vector operations. In all cases, integration follows the same ODE solver over the chosen $v$ while the $\mathcal{S}$ operator reconfigures bridges and directions on the fly.
4 GAF for Image Generation
4.1 Network Architecture
We present the GAF architecture and training procedure for image generation, training on both pixel-space and latent-space representations. The core components, trunk, twin predictors, and training objectives, remain unchanged across datasets; only the input/output dimensionality adapts to accommodate pixel- and latent-space inputs.
Figure 5: GAF architecture. The DiT-Trunk block follows the architecture of [williampeebles2023].
Trunk.
As shown in Figure 5, we employ a Diffusion Transformer (DiT) [williampeebles2023] architecture as our feature extractor. The GAF trunk takes in either pixel-space representation (e.g., $32\times 32\times 3$ px for low-resolution images) or a latent-space representation (e.g., $32\times 32\times 4$ for VAE, encoded from images with higher resolutions, such as $256\times 256\times 3$ px).
Patchify.
Patchify is the first layer of the trunk, converting the pixel- or latent-space inputs into a sequence of $T$ tokens. It does this by dividing the input in patches, and linearly embedding each patch into the trunk’s hidden dimension $D$. Standard sinusoidal positional embeddings are then added to the tokens to retain spatial information. As noted in [williampeebles2023], the choice of patch size $p$ presents a crucial tradeoff between computational cost and model granularity, as halving $p$ quadruples the GFLOPS with negligible effect on model parameter count. Therefore, we explore $p$ as $2$ and $4$ in our configurations.
Transformer Block.
The trunk processes the patch-embedded inputs through $N$ transformer blocks, each consisting of multi-head self-attention and feed-forward layers with adaptive layer normalization (adaLN). Conditioning on timestep $t$ and class label $y$ is achieved by embedding each into $D$-dimensional vectors. Timestep $t$ is mapped through sinusoidal encoding followed by a two-layer MLP. These embeddings are summed to form the conditioning vector $c=t_{emb}+y_{emb}$, which modulates each transformer block via adaLN. Following $N$ transformer blocks, a final adaLN layer conditioned on $c$ produces feature tokens $\mathbf{f}_{t}\in{\mathbb{R}^{N\times D}}$.
J Head (Noise Boundary.)
The $J$ head is a two-layer $\text{MLP}_{J}$, which maps trunk features to noise boundary residuals. The MLP expands from $D$ to $2D$ hidden units via a linear layer, applies GELU activation, and then projects to $p^{2}\times C$ features, where $C$ is the number of channels ($4$ for latents, 3 for RGB pixels). The final layer of $\text{MLP}_{J}$ is initialized to zero for stability. Finally, the outputs are reshaped (i.e., unpatchified) back to their spatial form ($C\times H\times W$).
K Heads (Data Boundaries).
For each class $k\in\{0,\dots,N-1\}$, each corresponding $K$ head is an independent two-layer $\text{MLP}_{K}$ with identical architecture to $J$. The $K$ heads share no parameters across classes, enabling independent transport operators for each class. During the forward pass, trunk features are computed once for the entire batch, then routed to the appropriate $K$ heads based on class labels. The routing ensures each $K$ head trains on only its assigned class, while the trunk learns shared representations across all classes. Notably, while we report results using this two-layer configuration, we empirically observed that a single linear projection is sufficient for high-fidelity generation, confirming that the backbone $\Phi$ captures the necessary feature information for the twins.
4.2 Training
We train GAF using three loss components as defined in (24) with mixed-precision training (bfloat16):
$\mathcal{L}_{\text{GAF}}=\mathcal{L}_{\text{pair}}+\lambda_{\text{res}}\,\mathcal{L}_{\text{res}}+\lambda_{\text{swap}}\,\mathcal{L}_{\text{swap}},$
where $\mathcal{L}_{\text{pair}}$ anchors the twins $J$ and $K$ to their endpoints ($\mathbf{z}_{y}$ and $\mathbf{z}_{x}$), $\mathcal{L}_{\text{res}}$ penalizes the residuals near endpoints, and $\mathcal{L}_{\text{swap}}$ enforces time-reversal symmetry at the midpoint for consistency and self-correction. Throughout all experiments, we set $\lambda_{\text{res}}=0.003$ and $\lambda_{\text{swap}}=0.002$ during training.
For each training sample, we draw a uniform timestep $t$ from $\text{Uniform}[t_{\text{eps}},1-t_{\text{eps}}]$, with $t_{\text{eps}}=1\times 10^{-3}$. We compute the forward prediction at $t$ for the main loss term, and evaluate it once more at the complementary timestep 1t1-t; this time-reversed prediction at 1t1-t is used to compute the swap loss.
4.3 Sampling
At inference, we generate samples by integrating the emergent velocity field v=KJv=K-J from noise ($t=0$) to data ($t=1$) using Euler’s method [10.5555/3600270.3602196]. Unlike standard models, GAF defines this field algebraically, enabling precise compositional generation directly within the ODE solver. As during training, we discretize the time interval $t_{\text{eps}},1-t_{\text{eps}}$] into uniformly spaced steps:
$t_{0}=t_{\text{eps}},t_{1},\dots,1-t_{\text{eps}},with~t_{\text{eps}}=1\times 10^{-3}$
At each step we update the latent state via
$z_{k+1}=z_{k}+(t_{k+1}-t_{k})v(z_{k},t_{k}).$ (29)
We initialize $\mathbf{z}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ at $t=t_{\text{eps}}$ and evolve to $t=1-t_{\text{eps}}$. For multi-class generation, samples are conditioned on their target class label $y$ through the corresponding $K_{n}$ heads.
For example, in the experiments in Section 5, we use $N=250$ Euler steps for high-quality evaluation. This translates to $250$ Number of Function Evaluations (NFE), as Eleur requires a single model evaluations per step. Additionally, to demonstrate the fast sampling capability, we also report results at $N=20$ (or NFE=$20$).
5 Experimental Evaluation
In this section, we evaluate GAF for image generation, as proposed in Section 4. We use $N=20$ Euler steps for all generations, with the exception of section 5.2, where we vary the number of steps.
5.1 Experimental Setup
5.1.1 Dataset
The proposed GAF model was trained on three datasets spanning different visual domains and classes.
CelebA-HQ [liu2015faceattributes] consists of $202,599$ celebrity face images resized to $64\times 64$ px, $256\times 256$ px, and $512\times 512$ px resolutions. We use this dataset for a single-class (i.e., one $J$ and one $K$) generation task.
AFHQ [choi2020starganv2] contains $15,000$ high-quality images of cats, dogs, and wildlife at $512\times 512$ px resolution. We use this dataset for a $3$-class (i.e., one $J$ and three $K$) generation task.
ImageNet-1k [deng2009imagenet, ILSVRC15] is the full ImageNet classification dataset with $1.28$ million training images across $1,000$ object classes, with $1,300$ images per classes. For demonstration purposes, we use two subsets: a single-class dataset (School Bus) and a multi-class subset consisting of $10$ distinct classes. For multi-class generation, we use $10$ $K$ heads, each corresponding to a specific class.
5.1.2 Training
We train CelebA-64 directly in pixel space, as the images are low-resolution. The other datasets contain higher-resolution images and are therefore first encoded by the SDXL-VAE [podell2024sdxl]. We use the AdamW optimizer with $\beta=(0.9,0.99)$, and learning rate $1\times 10^{-4}$ for all of the datasets, with no weight decay in any of the training runs. The ImageNet-1k bus class is trained for $300,000$ iterations, whereas the other models are trained for $500,000$ iterations.
5.2 Qualitative Evaluation
We qualitatively asses the effect of the number of Euler steps on generation quality using the CelebA-HQ dataset ($256\times 256$), utilizing the Fréchet Inception Distance (FID) [10.5555/3295222.3295408, Seitzer2020FID] as metric. All FID scores are computed using 50,000 generated samples. As shown in Figure 6, with very few steps (151-5) the samples are blurry but roughly coherent; increasing the steps to 104010-40 sharply improves the quality and semantic detail.
Figure 6: Effect of the number of Euler steps on sample quality. Columns correspond to $1,2,3,4$, $5,10,20,40,80,$ and $250$ Euler steps (left to right). Image sharpness and realism steadily improve as the number of steps increases.
The qualitative analysis shows a strong dependence on the number of Euler steps for CelebA-HQ 256 generation quality. The FID score experiences its most significant reduction initially, dropping sharply from $26.34$ at $20$ steps to $17.12$ at $40$ steps, corresponding to the rapid transition from blurry to sharp samples seen in Figure 7. Beyond $40$ steps, the curve exhibits clear diminishing returns: FID improves from about $12.024$ to only $10.30$ at $250$ steps. Therefore, while $250$ steps yields the best score, near-optimal, computationally efficient generation is already achieved around 8080-$100$ Euler steps.
Figure 7: Effect of the number of Euler steps on sample quality, measured by FID on CelebA-HQ $256$. FID decreases rapidly up to about $40$ steps, reflecting improved sample coherence, and then flattens, with strong diminishing returns beyond roughly $100$ steps.
Table 1: FID (50k samples) on CelebA-HQ-256 ($\times256$).
[…]
Glow [kingma_glow_2018] Flow 68.93
SDE [song2021scorebased] Score-based 7.23
LSGM [vahdat_score-based_2021] Latent score 7.22
LDM-4 [Rombach_2022_CVPR] Latent diffusion 5.11
GAF (ours) Endpoint regression 10.32
Table 1 reports FID scores on CelebA-HQ $256$ for a range of likelihood-based, score-based, and latent-diffusion models. GAF achieves a FID of $10.32$, placing it in a comparable range to established score-based methods on this dataset, while more heavily engineered latent-diffusion models still obtain the best reported numbers. This shows that our endpoint formulation is already competitive, with clear potential for further gains from architectural and training refinements.
5.3 Single-Class Generation (One $J$ and One $K$)
We first validate GAF on single-class generation, using the school bus class from ImageNet, which contains about 1.3k images used for training. Figure 8 shows some example generations. We can see that a single K head model with 250-step Heun sampling demonstrates high-quality, diverse generations. Note that the generations showcase varied viewpoints, lighting conditions, and backgrounds while maintaining class consistency.
Figure 8: Single-class generated samples ($256\times 256$ px) from GAF trained on the ImageNet school bus class.
5.4 Multi-Class Generation (One $J$ and Multiple $K$s)
We evaluate GAF on multi-class generation, using ImageNet-1k with $10$-class subsets and AFHQ ($512\times 512$ px) with three classes: cat, dog, and wild animals. Figure 9 and Figure 10 shows samples from each independent $K$ head, with each head corresponding to one of the ten or three classes with in each dataset. This demonstrates that GAF scales to multi-class generation while maintaining per-class quality. Each $K$ head produces diverse, high-fidelity samples within its class without interference from other heads.
Figure 9: Comprehensive gallery of generated samples from the $10$-class ImageNet subset. The figure demonstrates the model’s ability to handle high intra-class variance.
[…]
Figure 10: Multi-class generated samples ($512\times 512$ px) from GAF trained with one $J$ and three independent $K$ heads, on three classes (cat, dog, and wild) of the AFHQ dataset.
5.5 Transport Algebra
This section showcases three applications of GAF’s novel transport algebra: pairwise interpolation (in Section 5.5.1), continuous cyclic transport (in Section 5.5.2), and barycentric transport (in Section 5.5.3).
5.5.1 Pairwise Interpolation
We demonstrate direct $K$ head interpolation, by performing the operation $v=(1-\alpha)v_{i}+\alpha v_{j}$, for $\alpha\in[0,1].$ Figure 11 shows samples from interpolating between two classes from the AFHQ dataset, trained for multi-class generation in Section 5.4. We can see a smooth semantic transition between all three class pairs. Each interpolation produces coherent intermediate states, validating that independent $K$ heads enable linear composition in velocity space.
[…]
Figure 11: Pairwise interpolation samples ($512\times 512$ px) from GAF trained with one $J$ and three independent $K$ heads on the AFHQ dataset (i.e., the same model as sampled in Figure 10). Each row shows an interpolation along a different transport direction. Note that we did not explicitly train GAF for interpolation or translation. Instead, GAF enables such operations inherently using Transport Algebra.
5.5.2 Continuous Cyclic Transport
[…]
We observe that all 3000 samples achieve LPIPS=0.0. This indicates GAF’s transport algebra is deterministically reversible. The perfect closure validates that independent $K$ heads form a consistent algebraic structure where cyclic compositions return exactly to their starting point.
Figure 12 demonstrates continuous identity-preserving transport through multiple classes. Starting from an initial sample from the first class, we sequentially interpolate through two additional classes before returning to the original class, using velocity blending for each pairwise transition (as in Section 5.5.1). The endpoint of each interpolation becomes the starting point of the next, maintaining latent continuity throughout the cycle.
[…]
Figure 12: Continuous cyclic transport preserving identity across three classes. Each row shows a complete cycle starting from one class (left) to two interpolated classes, and finally returning to the initial class. Distinctive features (coloring, facial marking) are semantically preserved throughout each cycle while class-specific structure transforms.
5.5.3 Barycentric Transport
We extend pairwise interpolation to a three-way composition using barycentric coordinates. The blended velocity field is $v_{\text{blend}}=\alpha v_{i}+\beta v_{j}+\gamma v_{k}$, where $\alpha+\beta+\gamma=1$. We use square-to-simplex mapping with $\alpha=u(1-u),~\beta=(1-u)(1-v),$ and $\gamma=v$ to cover the full compositional space. Figure 13 shows three examples of three-way compositions, namely dog-cat-wild, cat-dog-wild, and wild-dog-cat. It shows all weighted combinations, with corners representing pure classes and the interior points showing multi-class blends.
(a) Dog-cat-wild.
(b) Cat-dog-wild.
(c) Wild-dog-cat.
vv01uu01Class 1 ($\beta=1$)Class 0 ($\alpha=1$)Class 2($\gamma=1$)
(d) Interpolation grid map showing the mixing ratios.
Figure 13: Barycentric Transport Algebra. Visualization of three-way velocity blending across the Dog, Cat, and Wild classes. (a-c) $7\times 7$ grids showing the semantic simplex. Corners represent pure classes; interior points show linear interpolations of the vector fields. (d) The corresponding weight map.
6 Conclusion
We introduced Generative Anchored Fields (GAF), a model that learns independent endpoint predictors $J$ and $K$ rather than a direct trajectory. The velocity field v=KJv=K-J emerges from the time-conditioned disagreement between these endpoints. This factorization enables Transport Algebra, allowing compositional control through vector arithmetic on learned $K$ heads. We demonstrated that GAF achieves strong sample quality while uniquely providing lossless cyclic transport (LPIPS=0.0) and precise semantic manipulation as intrinsic architectural primitives.
7 Future Work
GAF’s algebraic structure naturally extends to sequential domains. For video generation, we are reformulating Transport Algebra as Motion Algebra, where vector operations define transitions between frames. This will enable precise manipulation of temporal dynamics, offering a method to generate stable, controllable motion through the same arithmetic principles used for static data.
8 Acknowledgement
This work was funded in part by the Research Foundation Flanders (FWO) under Grant G0A2523N, IDLab (Ghent University-imec), Flanders Innovation and Entrepreneurship (VLAIO), and the European Union.