### v2

# Generative Anchored Fields: Controlled Data Generation via Emergent Velocity Fields and Transport Algebra

Abstract
We present Generative Anchored Fields (GAF), a generative model that learns independent endpoint predictors, $J$ (noise) and $K$ (data), from any point on a linear bridge. Unlike existing approaches that use a single trajectory or score predictor, GAF is trained to recover the bridge endpoints directly via coordinate learning. The velocity field v=KJv=K-J emerges from their time-conditioned disagreement. This factorization enables Transport Algebra: algebraic operations on multiple $J/K$ heads for compositional control. With class-specific $K_{n}$ heads, GAF defines directed transport maps between a shared base noise distribution and multiple data domains, allowing controllable interpolation, multi-class composition, and semantic editing. This is achieved either directly on the predicted data coordinates ($K$) using Iterative Endpoint Refinement (IER), a novel sampler that achieves high-quality generation in 585-8 steps, or on the emergent velocity field ($v$). We achieve strong sample quality (FID 7.51 on ImageNet $256\times 256$ and $7.27$ on CelebA-HQ $256\times 256$, without classifier-free guidance) while treating compositional generation as an architectural primitive. Code available at https://github.com/IDLabMedia/GAF.
1 Introduction
Modern generative models [16, 10, 13] achieve remarkable sample quality but lack precise compositional control, which is the ability to independently manipulate, interpolate, or combine learned class representations at inference time. Score-based diffusion and flow matching models learn a single time-conditioned predictor that maps noise to data via a unified field [38, 20]. Although models such as Stable Diffusion [32, 31] and FLUX.1 [3] generate images with high fidelity, their monolithic architecture treats control as an external steering process, implemented through classifier-free guidance [14], prompt engineering, latent space editing [26], or attention manipulation [11], rather than an intrinsic property of the learned representation.
Figure 1: Overview of GAF. (a) Samples from CelebA-HQ ($256\times 256$ px). (b) Multi-class composition with custom masks. (c) Barycentric interpolation: $v_{i\to j\to k}=\alpha v_{i}+\beta v_{j}+\gamma v_{k}$; corners are pure classes, interior points are weighted mixtures.
These limitations suggest a need to re-examine the core learning objectives. The trajectory-based paradigm focuses on learning step-by-step dynamics along a path [23]. This raises a fundamental question: must we learn a step-by-step mechanism of the trajectory or is it sufficient to know only its origin and destination? The trajectory-based paradigm assumes the first option. We explore the second one.
We present Generative Anchored Fields (GAF), where generation is an explicit algebraic operation. Rather than learning a single trajectory predictor, GAF learns independent endpoint operators $\{(J_{n},K_{n})\}_{n=1}^{N}$ that can be linearly combined to achieve precise compositional control of the generative process. Built on the principle of endpoint knowledge, the model uses a shared trunk, $\Phi$, which produces a time-conditioned feature bank to feed a pair of twin predictors, $J$ and $K$. The twins have opposing objectives: one is anchored to the noise distribution, the other to the data manifold. Unlike trajectory-based methods, GAF does not learn the velocity directly. Instead, the velocity field is an emergent consequence of the twins’ time-conditioned disagreement. Put simply, GAF is not trained to follow a path; it is trained only to know its endpoints.
This factorization of the transport operators ($J$ and $K$) has a direct consequence: precise and controllable generation. Because GAF learns independent endpoint predictors $\{(J_{n},K_{n})\}_{n=1}^{N}$, particularly when using class-specific $K_{n}$ heads, it naturally enables compositional operations at inference time. For example, one can switch between different $K$ heads mid-trajectory, interpolate between class manifolds with explicit geometric control, or arithmetically combine predictors to generate novel outputs. These operations are architectural primitives in GAF. We term this capability Transport Algebra: the ability to perform algebraic manipulations directly on the model’s learned components to achieve precise semantic control.
Our proposed model has the following contributions:
  • Coordinate-based Generative Modeling. Instead of learning a score, noise or velocity field, GAF learns the explicit coordinates of the transport endpoints: $J$ (noise) and $K$ (data). The velocity field $v(\mathbf{x}_{t},t)=K-J$ emerges as their antisymmetric disagreement. This formulation decouples the training objective (endpoint accuracy) from the sampler, allowing generation via Ordinary Differential Equation (ODE) integration or Iterative Endpoint Refinement (IER).
  • Transport Algebra. GAF introduces transport algebra to generative modeling: the ability to perform arithmetic operations directly on learned endpoint predictors. This enables controllable test-time generation, such as interpolation between classes, switching manifolds mid-trajectory, or combining multiple classes directly on the $K$ heads for novel compositional generation.
  • Iterative Endpoint Refinement. GAF has a native non-ODE sampler which iteratively refines endpoint estimates on a linear bridge, converging in 585-8 steps compared to 8025080-250 steps for Euler integration. IER forward uses only $K$ to refine the data endpoint, while IER reverse uses only $J$ to refine the noise endpoint, demonstrating that both the noise head $J$ and the data head $K$ are equally important.
[…]
  • Inherent modularity and extensibility. GAF’s shared trunk with pluggable “twin” or “tuplet” heads is inherently modular. One can easily add additional $K_{n}$ heads without altering the trunk.
  • Simple training and deterministic sampling. Training is a direct endpoint regression with simple residual and swap regularizers. Sampling is deterministic and guidance free, requiring small step Euler integration due to GAF’s self-correcting dynamics, unlike multi-step ODE solvers needed by other flow-based models or using IER.
2 Related Work
Modern Deep Generative Modeling (DGM) is largely dominated by methods that learn to reverse a data-to-noise trajectory, with key differences arising in the parametrization of this reverse process [43, 2, 20]. These approaches can be broadly categorized into score-based, denoising and velocity-based formulations. Our work departs from these formulations by introducing an endpoint regression objective for underlying transport dynamics.
Score-based generative models learn the score function (i.e., the gradient of the log-density $\nabla_{\mathbf{x}}\log p(\mathbf{x}_{t})$) of noise-perturbed data distribution [41, 42]. By training a noise conditional score network across multiple noise scales, these models can generate samples via annealed Langevin dynamics. This paradigm was later formalized through the lens of stochastic differential equations (SDE), which established the score function as the key component needed to reverse a continuous-time forward diffusion process [43]. This SDE framework provided a unifying mathematical language for a wide variety of diffusion models, but the core objective remained the estimation of the score.
Concurrently, Denoising Diffusion Probabilistic Models (DDPMs) [38, 13] achieved state-of-the-art sample quality by simplifying the transport objective to predict the noise $\epsilon$ added at each step of a discrete forward process. This noise-prediction ($\epsilon$-prediction) target proved highly stable and effective, becoming the default standard, such as in OpenAI’s DALL-E [7], Google’s Imagen [34], and Midjourney [28]. Later, Denoising Diffusion Implicit Models (DDIM) [39] introduced deterministic sampling for diffusion models to remove randomness during generation, allowing the model to skip steps and produce high-quality images. Subsequent improvements focused on optimization, the training recipe, and architecture, with new methods like cosine noise schedules, learned variances [27], architectural upgrades [7], and powerful conditioning mechanisms [14]. A pivotal development was classifier-free guidance (CFG) [14], which enabled high-fidelity conditional generation from a single model by extrapolating between conditional and unconditional outputs. While our work uses the powerful transformer backbone (DiT) [45, 8, 30] that pioneered in this domain, we depart from the standard practice of predicting either the score or the noise.
Most recently, the field has increasingly shifted towards a deterministic, continuous-time perspective rooted in ODEs and optimal transport [46, 20, 23, 21]. These models learn a velocity field $v_{\theta}(\mathbf{x}_{t},t)$ that governs a probability flow ODE, transporting samples from a simple base distribution (such as a Gaussian distribution) to the data distribution. While mathematically related to the score, this velocity ($v$-prediction) parametrization was shown to have better numerical stability and scaling properties [36, 15]. The development of simulation-free training objectives, such as Flow Matching [20] and Rectified Flows [23], has made this approach highly efficient. Stochastic Interpolants [1] generalize this framework by constructing interpolating paths between arbitrary distributions, sharing GAF’s bridge formulation but learning a velocity field rather than endpoint coordinates. Consistency Models [40] target few-step generation by distilling a pretrained diffusion model into a single-step predictor; GAF’s IER achieves comparable step efficiency (585-8 steps) without a pretrained teacher or distillation stage. By directly regressing the network’s output onto a target field, often defining a straight path between noise and data, these methods learn linear trajectories, enabling high-quality generation in very few sampling steps, as demonstrated by models like FLUX.1 [3], Stable Diffusion 3 [9], and InstaFlow [24].
All these approaches, whether predicting score, noise, or velocity, learn a single time-conditioned predictor. While effective for generation, this monolithic architecture limits compositional control: one cannot independently manipulate learned representations for different classes or combine them arithmetically. Conditional generation instead relies on guidance mechanisms like CFG or prompt engineering. Recent work has explored alternative control mechanism. For example, Compositional Visual Generation [22] defines two compositional operators (AND and NOT) for multi-concept generation, using a set of diffusion models to enabling logical operation over concepts. Moreover, ControlNet [47] enables spatial control in text-to-image generation by locking the original diffusion model while training another network for conditional control with extra inputs like edges or depth. Although these methods provide impressive control, they require either maintaining multiple models for composing their outputs or training additional conditioning networks. Both approaches treat control as an add-on mechanism rather than an intrinsic architectural property.
To address this limitation, GAF introduces an algebraic factorization of the transport dynamics. Instead of learning a single, monolithic velocity field $v_{\theta}$, GAF learns two endpoint predictors: a twin $J$ that regresses towards the noise distribution and a twin $K$ that regresses towards the data manifold. The velocity field $v$ then emerges from their difference, v=KJv=K-J, enforced by a paired and time-antisymmetric loss that promotes consistency between the forward and backward (reverse in time) dynamics. This explicit factorization of the transport operators ($J$ and $K$) enables unique architectural choices, such as independent, class-specialized $K$-heads. Unlike external composition [22] or auxiliary networks [47], GAF’s compositional capabilities are intrinsic: they emerge from the factorized architecture. This design unlocks novel inference-time capabilities, allowing vector operations between class distributions, a process we term Transport Algebra. This form of algebraic manipulation over $\{(J_{n},K_{n})\}_{n=1}^{N}$ to generate new samples with precise compositional control directly is not naturally afforded by standard score-, flow-, or noise-based conditional generative models.
A summary of how GAF compares to dominant trajectory-based paradigms is provided in Table 1.
Table 1: A comparison of GAF with Score-Based Diffusion and Flow Matching models across key aspects.
Feature Score-Based Diffusion Flow Matching GAF
Foundation Stochastic (SDE) Deterministic (ODE) Deterministic (IER/ODE)
Primary Learning Target A Score Function $\nabla_{x}\log{p}_{\theta}(x)$ Velocity Field ($v$) Coordinate Learning: Endpoints via twins, $J\!\to\!\mathbf{z}_{y},\ K\!\to\!\mathbf{z}_{x}$
Nature of Dynamics Derived (Indirect) Prescribed (Direct) Direct (IER); Emergent ($v$)
Velocity’s Role Derived from the score function. The direct regression target. A byproduct of endpoints (KJK-J)
Model’s Task Denoising/Score estimation Velocity Regression Endpoint Regression
3 Generative Anchored Fields (GAF)
GAF is a generative model that learns independent endpoint predictors rather than a trajectory or velocity field. In Section 3.1, we define the bridge formulation, the twin predictors $J$ and $K$, and derive the training objective. In Section 3.2, we discuss the structural properties that follow from this factorization. In Section 3.3, we introduce Transport Algebra, the compositional operations enabled by independent endpoint heads.
3.1 The GAF Model
We introduce the Generative Anchored Fields model. Let $\mathbf{z}_{x},\mathbf{z}_{y}\in\mathbb{R}^{d}$ be samples from a data distribution $\mathbf{z}_{x}\sim p_{\text{data}}(\mathbf{x})$ and a standard normal distribution $\mathbf{z}_{y}\sim\mathcal{N}(0,I)$, respectively. Additionally, let $\mathbf{c}\in{\mathbb{R}^{d}}$ be an optional conditioning vector (e.g., a class embedding).
The model operates on linear bridges connecting these samples:
$\mathbf{x}_{t}=(1-t)\mathbf{z}_{y}+t\mathbf{z}_{x},\quad t\in[0,1].$ (1)
The central task is for GAF to learn the two anchor endpoints $\mathbf{z}_{y}$ and $\mathbf{z}_{x}$ from a given point on the bridge that they define. We call this task Coordinate Learning: the ability to identify the coordinates (endpoints) of a bridge from any point along its trajectory. The core of the model is a neural network, referred to as the trunk $\Phi$, that processes the bridge $\mathbf{x}_{t}$ and timestep $t$ to produce a feature bank $\mathbf{f}_{t}$:
$\mathbf{f}_{t}=\Phi(\mathbf{x}_{t},t).$ (2)
Optionally, for class guidance we condition on $\mathbf{c}$:
$\mathbf{f}_{t}=\Phi(\mathbf{x}_{t},t,\mathbf{c}).$ (3)
The feature bank $\mathbf{f}_{t}$ is then processed by a pair of twins $J$ and $K$, which regress the bridge $\mathbf{x}_{t}$ towards its endpoints (i.e., noise $\mathbf{z}_{y}$ and data $\mathbf{z}_{x}$, respectively). The twins are formally defined as:
$J$ $:=(1-t)\mathbf{x}_{t}+H_{J}(\mathbf{f}_{t}),$ (4)
$K$ $:=t\mathbf{x}_{t}+H_{K}(\mathbf{f}_{t}).$ (5)
$H_{J}$ and $H_{K}$ are identical neural network architectures, but trained for separate tasks. We denote $J_{\text{res}}:=H_{J}(\mathbf{f}_{t})$ and $K_{\text{res}}:=H_{K}(\mathbf{f}_{t})$. We term our model “Anchored Fields” because the predictors $J$ and $K$ are anchored to fixed endpoints: $J$ targets the noise distribution, $K$ targets the data manifold. Unlike trajectory-based models that learn the dynamics between these points, GAF anchors its learning directly to the endpoints themselves.
3.1.1 Emergent Transport Dynamics
The twins move in opposite directions along the bridge. More specifically, $J$ is oriented toward the noise endpoint, and $K$ toward the data endpoint. Their ideal boundary conditions are $J=\mathbf{z}_{y}$ at $t=0$, and $K=\mathbf{z}_{x}$ at $t=1$. To quantify the twins’ motion relative to each other along the bridge, we compute the path’s instantaneous velocity with respect to $t$. For the linear bridge (1), we have:
$\frac{d\mathbf{x}_{t}}{dt}=\frac{d}{dt}\big[(1-t)\mathbf{z}_{y}+t\mathbf{z}_{x}\big]=\mathbf{z}_{x}-\mathbf{z}_{y}.$ (6)
With this ground-truth velocity in hand, we define the learned velocity field $v$ as the twins disagreement:
$v(\mathbf{x}_{t},t)=K-J.$ (7)
This learned field then specifies an ODE that transports the noise sample $\mathbf{z}_{y}$ to the data sample $\mathbf{z}_{x}$:
$\frac{d\mathbf{{x}}_{t}}{dt}=v(\mathbf{x}_{t},t).$ (8)
We term $J$ and $K$ twins to emphasis their symmetric roles: both are endpoint predictors that must be learned in balance. Their velocity field v=KJv=K-J emerges from their difference, so imbalanced training would bias their dynamics towards one endpoint.
We generate new data samples by either (i) iteratively refining the endpoints via the IER algorithm or (ii) integrating the velocity field ODE from $t=0$ to $t=1$ along a linear or cosine time grid, using Euler integration [4, 15] (see Section 4.4).
3.1.2 GAF Training
The model is trained using the twins pair loss:
$\mathcal{L_{\text{pair}}}=\mathbb{E}_{\mathbf{z}_{x},\mathbf{z}_{y},t}[(1-t)\|J-\mathbf{z}_{y}\|_{2}^{2}+t\|K-\mathbf{z}_{x}\|_{2}^{2}]$ (9)
Residual penalty. To improve training stability, we also enforce the model’s boundary behavior. That is, we examine the prediction errors at the ideal boundary condition $J-\mathbf{z}_{y}$ and $K-\mathbf{z}_{x}$ by expanding them at the endpoints.
From $J=(1-t)\mathbf{x}_{t}+J_{\text{res}}$, at $t=0$, we have $J=\mathbf{z}_{y}+J_{\text{res}}$. Thus, achieving the ideal condition $J=\mathbf{z}_{y}$ requires $J_{\text{res}}=0$. By substituting the definition of $J$, we get:
$J=\mathbf{z}_{y}\equiv 0$ $=J-\mathbf{z}_{y}$ (10)
$=(1-t)\mathbf{x}_{t}+J_{\text{res}}-\mathbf{z}_{y}$
$=(1-t)((1-t)\mathbf{z}_{y}+t\mathbf{z}_{x})+J_{\text{res}}-\mathbf{z}_{y}$
$=((1-t)\mathbf{z}_{y}+t\mathbf{z}_{x})-(t(1-t)\mathbf{z}_{y}-t^{2}\mathbf{z}_{x})+J_{\text{res}}-\mathbf{z}_{y}$
$=(1-t)\mathbf{z}_{y}+t\mathbf{z}_{x}-t^{2}\mathbf{z}_{x}-t(1-t)\mathbf{z}_{y}+J_{\text{res}}-\mathbf{z}_{y}$
$\equiv\underbrace{t(1-t)(\mathbf{z}_{x}-\mathbf{z}_{y})}_{\text{cross term}}-\underbrace{t\mathbf{z}_{y}}_{\text{endpoint}}+\underbrace{J_{\text{res}}}_{\text{residual}}$ (11)
Since our ideal learning goal is to have $J=\mathbf{z}_{y}$ at $t=0$, the ideal boundary condition implies the magnitude of the learned residual term $J_{\text{res}}$ must be zero at this point: $J_{\text{res}}|_{t=0}=0$.
Similarly, from $K=t\mathbf{x}_{t}+K_{\text{res}}$, at $t=1$, $K=\mathbf{z}_{x}+K_{\text{res}}$. Thus, achieving the ideal condition $K=\mathbf{z}_{x}$ requires $K_{\text{res}}=0$. By substituting the definition of $K$, we get:
$K=\mathbf{z}_{x}\equiv 0$ $=K-\mathbf{z}_{x}$ (12)
$=t\mathbf{x}_{t}+K_{\text{res}}-\mathbf{z}_{x}$
$=t((1-t)\mathbf{z}_{y}+t\mathbf{z}_{x})+K_{\text{res}}-\mathbf{z}_{x}$
$=t(1-t)\mathbf{z}_{y}+t^{2}\mathbf{z}_{x}-\mathbf{z}_{x}+K_{\text{res}}$
$=t(1-t)\mathbf{z}_{y}+\mathbf{z}_{x}(t^{2}-1)+K_{\text{res}},\quad t^{2}-1=(1-t)(-1-t)$
$=t(1-t)\mathbf{z}_{y}+\mathbf{z}_{x}(1-t)(-1-t)+K_{\text{res}}$
$=(1-t)(t\mathbf{z}_{y}+\mathbf{z}_{x}(-1-t))+K_{\text{res}}$
$=(1-t)(t\mathbf{z}_{y}-\mathbf{z}_{x}-t\mathbf{z}_{x})+K_{\text{res}}$
$=(1-t)(t(\mathbf{z}_{y}-\mathbf{z}_{x})-\mathbf{z}_{x})+K_{\text{res}}$
$=t(1-t)(\mathbf{z}_{y}-\mathbf{z}_{x})-(1-t)\mathbf{z}_{x}+K_{\text{res}}$
$\equiv\underbrace{t(1-t)(\mathbf{z}_{y}-\mathbf{z}_{x})}_{\text{cross term}}-\underbrace{(1-t)\mathbf{z}_{x}}_{\text{endpoint}}+\underbrace{K_{\text{res}}}_{\text{residual}}$ (13)
Since our ideal learning goal is to have $K=\mathbf{z}_{x}$ at $t=1$, the ideal boundary condition implies the magnitude of the learned residual term $K_{\text{res}}$ must be zero at this point: $K_{\text{res}}|_{t=1}=0$.
In both derivations in (11) and (13), the cross term represents the interaction between the endpoints. The factor t(1t)t(1-t) acts as a symmetric weighting or modulation factor. Its weight peaks at the midpoint, $t=\frac{1}{2}$, concentrating the error signal at this point of maximum complexity. This forces both $J$ and $K$ to prioritize learning the correct transition path, rather than just matching the endpoints themselves.
While the primary endpoint loss (9) implicitly encourages the ideal boundary conditions, we introduce an explicit residual regularization penalty to enforce them directly and improve stability. From (11) and (13), we have our residual penalty as:
$\mathcal{L_{\text{res}}}=(1-t)\|J_{\text{res}}\|_{2}^{2}+t\|K_{\text{res}}\|_{2}^{2}$ (14)
Residual-antisymmetry penalty. The endpoint residual penalty ensures $J\rightarrow\mathbf{z}_{y}$ as $t\rightarrow 0$ and $K\rightarrow\mathbf{z}_{x}$ as $t\rightarrow 1$, but it does not fully determine the intermediate behavior of GAF on the bridge. Since we want GAF to behave consistently across the entire bridge, we need a mechanism to do that. One way to do that is by swapping the residuals $J_{\text{res}}$ and $K_{\text{res}}$ across the bridge and create antisymmetric relationship between them. We define the swap operator (function) acting on the residuals at $t$ and 1t1-t as:
$\mathcal{S}:(\mathbf{z}_{x},\mathbf{z}_{y},t,c)\mapsto(\mathbf{z}_{y},\mathbf{z}_{x},1-t,c)$ (15)
Using our bridge formalism in (1) and the swap operator in (15), we have:
$\mathbf{x}_{1-t}$ $=(1-(1-t))\mathbf{z}_{x}+(1-t)\mathbf{z}_{y}$ (16)
$\mathbf{x}_{1-t}$ $=t\mathbf{z}_{x}+(1-t)\mathbf{z}_{y}$ (17)
From (1) and (17), $\mathbf{x}_{t}\stackrel{{\scriptstyle\mathcal{S}}}{{=}}\mathbf{x}_{1-t}$, i.e., the swap stays on the bridge, and only the direction of travel changes.
Here, $\stackrel{{\scriptstyle\mathcal{S}}}{{=}}$ means equality after applying the swap (namely, swapping endpoints and flipping time).
Using our trunk from (3) and swap function from (15), the endpoint-swap antisymmetric residuals are defined as:
$\tilde{J}_{\text{res}}:=H_{J}(\Phi(\mathbf{x}_{1-t},1-t,c)),\quad\tilde{K}_{\text{res}}:=H_{K}(\Phi(\mathbf{x}_{1-t},1-t,c)).$ (18)
Similarly, we have:
$v(t)$ $=t\mathbf{x}_{t}+K_{\text{res}}-((1-t)\mathbf{x}_{t}+J_{\text{res}})$ $v(1-t)$ $=(1-t)\mathbf{x}_{1-t}+\tilde{K}_{\text{res}}-((1-(1-t))\mathbf{x}_{1-t}+\tilde{J}_{\text{res}})$
$=(2t-1)\mathbf{x}_{t}+(K_{\text{res}}-J_{\text{res}})$ $=(1-2t)\mathbf{x}_{1-t}+(\tilde{K}_{\text{res}}-\tilde{J}_{\text{res}})$ (19)
Our residual antisymmetric targets are:
$J_{\text{res}}=-\tilde{K}_{\text{res}},\quad K_{\text{res}}=-\tilde{J}_{\text{res}}\;\Rightarrow\;v(1-t)=-v(t)$ (20)
Thus, the swap function flips the velocity field in time.
Although GAF is trained with an endpoint regression, its emergent velocity field exhibits how it learns the transport dynamics. From (19), we have $v(\mathbf{x}_{t},t)=(2t-1)\mathbf{x}_{t}+(K_{\text{res}}-J_{\text{res}})$. Let $\Delta_{\text{res}}=K_{\text{res}}-J_{\text{res}}$ and examine $v$ at $t\in[0,1]$.
$v(\mathbf{x}_{t},t)=\begin{cases}-\mathbf{z}_{y}+\Delta_{\text{res}}&\text{lim}_{t\rightarrow 0}\\ \Delta_{\text{res}}&t=\frac{1}{2}\\ \mathbf{z}_{x}+\Delta_{\text{res}}&\text{lim}_{t\rightarrow 1}\end{cases}$ (21)
Similarly, from the swap operator in (19), we have $v(\mathbf{x}_{1-t},t)=(1-2t)\mathbf{x}_{1-t}+(\tilde{K}_{\text{res}}-\tilde{J}_{\text{res}})$. Let $\tilde{\Delta}_{\text{res}}=\tilde{K}_{\text{res}}-\tilde{J}_{\text{res}}$ and examine $v$ at $t\in[0,1]$.
$v(\mathbf{x}_{1-t},t)=\begin{cases}\mathbf{z}_{y}+\tilde{\Delta}_{\text{res}}&\text{lim}_{t\rightarrow 0}\\ \tilde{\Delta}_{\text{res}}&t=\frac{1}{2}\\ -\mathbf{z}_{x}+\tilde{\Delta}_{\text{res}}&\text{lim}_{t\rightarrow 1}\end{cases}$ (22)
As we have discussed earlier, in GAF, we do not train the velocity field directly, as in rectified flow (see Section 5.7 for a comparison). Instead, the velocity simply emerges from the twins opposite movement. However, from (21) and (22), we can see that the emergent velocity field incorporates knowledge of the trajectory’s endpoint.
Let $g_{0}=J_{\text{res}}+\tilde{K}_{\text{res}}$ and $g_{1}=K_{\text{res}}+\tilde{J}_{\text{res}}$. The ideal time-antisymmetric condition is $g_{0}=g_{1}=0$. Hence, we define the residual-antisymmetry penalty as:
$\mathcal{L}_{\text{swap}}$ $=\|g_{0}\|_{2}^{2}+\|g_{1}\|_{2}^{2}$
$=\|J_{\text{res}}+\tilde{K}_{\text{res}}\|_{2}^{2}+\|K_{\text{res}}+\tilde{J}_{\text{res}}\|_{2}^{2}.$ (23)
From (9), (14), and (23), the GAF training loss is
$\mathcal{L}_{\text{GAF}}=\mathcal{L}_{\text{pair}}+\lambda_{\text{res}}\,\mathcal{L}_{\text{res}}+\lambda_{\text{swap}}\,\mathcal{L}_{\text{swap}}.$ (24)
where $\lambda_{\text{res}}$ and $\lambda_{\text{swap}}$ are hyperparameters (e.g, both $0.01$).
The overall architecture of GAF and its data flow is summarized in Figure 2, illustrating how the training components (top) predict endpoints that drive sampling strategies (bottom) via IER or velocity integration.
Figure 2: The high-level overview of GAF architecture and its data flow.
3.2 Properties of GAF
3.2.1 Main Components
GAF comprises four important components. Each component is necessary for the model to retain its defining behavior; removing one of them either destroys the emergent velocity, or reduces GAF to a different formulation or eliminates learning altogether.
  1. 1.
    Trunk (shared backbone). The trunk $\Phi(\mathbf{x}_{t},t)\in\mathbb{R}^{d}$ provides the shared time-conditioned representation used by the twins. It hides architectural complexity details from the twins and enables modularity (plugging different heads, conditioning, or label embeddings) and parameter sharing. Removing the trunk would destroy this modularization ability and prevent the twins from coordinating through a common state. Hence, if you remove the trunk, you would lose the modularity of GAF.
  2. 2.
    Twins (opposing heads). The two time-conditioned predictors $J$, and $K$ pull the anchored state $\mathbf{x}_{t}$ toward noise and data. Their antisymmetric disagreement $v(\mathbf{x}_{t},t)=K-J$ yields the emergent velocity field. Removing either head eliminates the disagreement, the emergent velocity, and, by design, GAF itself.
  3. 3.
    Twin Pair Loss (endpoint anchoring). Each twin is trained with its endpoint reconstruction. The pairs are required to anchor both ends of the path. Dropping one of the losses collapses GAF.
  4. 4.
    Sampler. GAF samples natively via Iterative Endpoint Refinement (IER), which refines endpoint estimates direclty on the bridge withouth solving an ODE. Alternatively the emergent field v=KJv=K-J can be integrated with a single ODE solver (e.g., Euler or Heun). In both cases, no auxiliary corrector network, guidance model, or second model is required.
3.2.2 GAF is Naturally Modular
GAF is a naturally modular model. This means that one can simply add additional $J$ and $K$ heads on to the trunk to add new members to the twins or “tuplets”, without changing any other part of the architecture. Each head represents a different noise region and modality. For example, one could simultaneously learn an image with its albedo ($K_{1}$), depth map ($K_{2}$), and normal map ($K_{3}$), with each modality learned by specific member of the twins. We can add tuplets ($K_{n}$) as long as the base trunk is capable of feeding enough features for all tuplets, and as long as it is designed for the specific modality.
GAF is modal-agnostic. That is, to change the modality of GAF, we only need to change the trunk’s modality and the I/O. For example, to change the modality from image data to video data, you only need to change the trunk. The other GAF components need no (or minimal) change.
3.3 Transport Algebra
Transport Algebra organizes GAF’s endpoint pairs into an algebra of endpoint predictors, enabling composable operations that yield controlled data generation. Recall that GAF defines two time-conditioned endpoints $J$ and $K$. Because GAF is modular, the shared trunk can host multiple tuplets $\{(J_{n},K_{n})\}_{n=1}^{N}$. This creates a natural algebra in which compositional operations can be performed (i) directly on the endpoint predictors by combining $J_{n}$ and $K_{n}$ and then applying endpoint refinement, and (ii) by using the emergent velocity field $v_{n}=K_{n}-J_{n}$ and combining these fields algebraically. This ability of GAF to operate algebraically leads to a precise and semantically coherent method of data generation across multiple classes and domains.
Bridge and swap operators.
The swap operator from (15) carves a two-way lane on a bridge and enforces a bidirectional highway for transport on the same bridge. Using the swap operator, we define three elementary operators on the bridge configurations for multiple path traversals as follows:
$\mathcal{S}_{\text{swap}}$ $:(\mathbf{z}_{y},\mathbf{z}_{x},t,c)\mapsto(\mathbf{z}_{x},\mathbf{z}_{y},t,c)$ (Swap endpoints) (25)
$\mathcal{S}_{\text{flip}}$ $:(\mathbf{z}_{y},\mathbf{z}_{x},t,c)\mapsto(\mathbf{z}_{y},\mathbf{z}_{x},1-t,c)$ (Flip time) (26)
$\mathcal{S}_{\text{swap \& flip}}$ $:(\mathbf{z}_{y},\mathbf{z}_{x},t,c)\mapsto(\mathbf{z}_{x},\mathbf{z}_{y},1-t,c)$ (Swap & flip) (27)
From (25) and (26), we get $\mathbf{x}_{t}=(1-t)\mathbf{z}_{x}+t\mathbf{z}_{y}$. Just like the main bridge in (15), the new swaps in (25) and (26) likewise remain on a bridge. Consequently, one can compose these operators to perform controlled generation at test time, using the resulting configurations to traverse or switch modalities, see Section 5.6. Figure 3 illustrates the four swap configurations.
Figure 3: Visualization of the swap operator $\mathcal{S}$ acting on a bridge. (A) The initial configuration of the bridge (1). Subfigures (B), (C), and (D) illustrate the results of applying the operations described in equations (25), (26), and (27), respectively.
$J/K$ pairing topologies.
The swap operator $\mathcal{S}$ gives us the mechanism to traverse through the endpoints back and forth in time. In addition, we can also consider the three main different ways multiple pairs of twin $J$ and $K$ can be mapped into a $\{(J_{n},K_{n})\}_{n=1}^{N}$ mappings. The pairings are: (A) one-to-one, (B) star (one-to-many), and (C) clustered mapping, shown in Figure 4.
Figure 4: Three $J/K$ pairing topologies: (A) one-to-one, (B) star (one-to-many), and (C) clustered. Each topology shows different ways to map the noise endpoint $J$ to the data endpoint $K$.
Transport algebra operations.
GAF’s independent $K$ heads enable compositional transport through endpoint operations (IER) and through linear velocity operations. Let $i,j\in\{1,\dots,N\}$.
  1. 1.
    Single-class transport. Samples from class $i$ can be generated by applying IER algorithm using the corresponding $K_{i}$ head to iteratively refine the data endpoint from $\mathbf{z}_{y}\sim\mathcal{N}(0,I)$.
    If velocity view is used and the noise endpoint $J$ is shared, then
    $v_{i}=K_{i}-J$
    If each modality has its own noise endpoint $J_{i}$, then:
    $v_{i}=K_{i}-J_{i}$
    Integrating $v_{i}$ from noise ($t=0$) to data ($t=1$) generates samples from class $i$.
  2. 2.
    Cross-modality transport (via $J$).
    Move between data endpoints of two (or more) domains. To move from class $i$ to class $j$, we use $J$ as a terminal and $K_{i\to j}$$K_{i}\xrightarrow{J}K_{j}$ (IER) or $v_{i\to j}=v_{i}\xrightarrow{-v_{i}}J\to v_{j}$ (ODE). The $v_{i\to j}$ expands to an encode-decode process:
    1. (a)
      Decode to class $i$: Integrate forward with $v_{J\to i}=K_{i}-J$ to reach $K_{i}$ (forward time)
    2. (b)
      Encode to noise: Integrate with $v_{i\to J}=J-K_{i}$ to reach $J$ (reverse time)
    3. (c)
      Decode to target class $j$: Integrate forward with $v_{J\to j}=K_{j}-J$ to reach $K_{j}$ (forward time)
      The full transport path is:
      $v_{i\to j}=K_{i}\xrightarrow{-(K_{i}-J)}J\xrightarrow{K_{j}-J}K_{j}$
    This sequential transport process uses $J$ as a universal intermediate representation shared across all classes, enabling transport and composition between any pair of $K$ heads.
  3. 3.
    Direct $K$-head interpolation. We can interpolate directly on $K$ heads or in the velocity space via weighted combinations
    $K_{\alpha}=(1-\alpha)K_{i}+\alpha K_{j},$ (28)
    $v_{\alpha}=(1-\alpha)(K_{i}-J)+\alpha(K_{j}-J)=(1-\alpha)v_{i}+\alpha v_{j},$ (29)
    for $\alpha\in[0,1]$.
    A simultaneous weighted combination of multiple $K$ heads enables compositional generation, which is experimented on in Section 5.6.3. These operations demonstrate GAF’s transport algebra where $K$ independent heads act as linear operators that can be composed, interpolated, and combined to traverse the generative space. This property extends naturally to sequential domains. For example, for video generation, we hypothesize that GAF can smoothly transition between temporal states while preserving semantic coherence through IER interpolation or velocity interpolation, which we will explore in future work.
  4. 4.
    Multi-class composition (barycentric). Combine multiple $K$ heads or velocity fields by linear vector operations.
    $K_{\text{blend}}=\sum_{m}w_{m}K_{m}$ (30)
    $v_{\text{blend}}=\sum_{m}w_{m}(K_{m}-J)$ (31)
    where $\sum_{m}w_{m}=1$
By using $J_{n},K_{n}$, IER, $v_{n}$ and the $\{(J_{n},K_{n})\}_{n=1}^{N}$ pairings, together with the swap operator $\mathcal{S}$ (which swaps $J\leftrightarrow K$, $z_{y}\leftrightarrow z_{x}$, and $t\leftrightarrow 1-t$) one can traverse modalities, switch modalities mid generation, revert to the previous modality, chain back to the origin, or even use a specific generated modality as a bootstrap for a chained generation. All of this is possible by switching endpoint heads and applying endpoint refinement, or by switching between the emergent velocity field $v_{i}$, $i\in\{1,\dots,N\}$, and applying simple vector operations. In all cases, $\mathcal{S}$ operator reconfigures bridges and directions on the fly.
4 GAF for Image Generation
We present the GAF architecture and training procedure for image generation, training on both pixel-space and latent-space representations. The core components, trunk, twin predictors, and training objectives, remain unchanged across datasets; only the input/output dimensionality adapts to accommodate pixel- and latent-space inputs.
4.1 Network Architecture
An overview of the architecture for image generation is given in Figure 5. We discuss each component below.
Figure 5: GAF architecture for image generation. The DiT-Trunk block follows the architecture of [30]. GAF replaces DiT’s single output projection with two independent heads, $H_{J}$ and $H_{K}$ (see Section 3.1), each producing endpoint residuals that are unpatchified separately and combined with scaled input to form the twins $J$ and $K$.
Trunk.
As shown in Figure 5, we employ a Diffusion Transformer (DiT) [30] architecture as our feature extractor. The GAF trunk takes in either pixel-space representation (e.g., $32\times 32\times 3$ px for low-resolution images) or a latent-space representation (e.g., $32\times 32\times 4$, encoded from high-resolution images such as $256\times 256\times 3$ px using a pretrained variational autoencoder (VAE) [16]). That is because, at higher resolution, pixel-space processing becomes impractical due to memory and compute constraints.
Patchify.
Patchify is the first layer of the trunk, converting the pixel- or latent-space inputs into a sequence of $T$ tokens. It does this by dividing the input in patches, and linearly embedding each patch into the trunk’s hidden dimension $D$. Standard sinusoidal positional embeddings are then added to the tokens to retain spatial information. As noted in [30], the choice of patch size $p$ presents a crucial tradeoff between computational cost and model granularity, as halving $p$ quadruples the GFLOPS with negligible effect on model parameter count. Therefore, we set $p$ as $2$ in our configurations.
DiT Block.
The trunk processes the patch-embedded inputs through $N$ transformer blocks, each consisting of multi-head self-attention and feed-forward layers with adaptive layer normalization (adaLN). Conditioning on timestep $t$ and class label $y$ is achieved by embedding each into $D$-dimensional vectors. Timestep $t$ is mapped through sinusoidal encoding followed by a two-layer MLP. These embeddings are summed to form the conditioning vector $c=t_{emb}+y_{emb}$, which modulates each transformer block via adaLN. Following $N$ transformer blocks, a final adaLN layer conditioned on $c$ produces feature tokens $\mathbf{f}_{t}\in{\mathbb{R}^{N\times D}}$.
J Head (Noise Boundary.)
The $J$ head is a single-layer $\text{MLP}_{J}$, which maps trunk features to noise boundary residuals. The MLP expands from $D$ to $2D$ hidden units via a linear layer, and then projects to $p^{2}\times C$ features, where $C$ is the number of channels ($4$ for latents, 3 for RGB pixels). The final layer of $\text{MLP}_{J}$ is initialized to zero for stability. Finally, the outputs are reshaped (i.e., unpatchified) back to their spatial form ($C\times H\times W$).
K Heads (Data Boundaries).
For each class $k\in\{0,\dots,N-1\}$, each corresponding $K$ head is an independent single-layer $\text{MLP}_{K}$ with identical architecture to $J$. The $K$ heads share no parameters across classes, enabling independent transport operators for each class. During the forward pass, trunk features are computed once for the entire batch, then routed to the appropriate $K$ heads based on class labels. The routing ensures each $K$ head trains on only its assigned class, while the trunk learns shared representations across all classes. Therefore, a single linear projection is sufficient for high-fidelity generation, confirming that the backbone $\Phi$ captures the necessary feature information for the twins.
4.2 Training
We train GAF with three loss components defined in (24) using mixed-precision (bfloat16) training.
$\mathcal{L}_{\text{GAF}}=\mathcal{L}_{\text{pair}}+\lambda_{\text{res}}\,\mathcal{L}_{\text{res}}+\lambda_{\text{swap}}\,\mathcal{L}_{\text{swap}},$
where $\mathcal{L}_{\text{pair}}$ anchors the twins $J$ and $K$ to their endpoints ($\mathbf{z}_{y}$ and $\mathbf{z}_{x}$), $\mathcal{L}_{\text{res}}$ penalizes the residuals near endpoints (as $t\rightarrow 0$ for $J$ and as $t\rightarrow 1$ for $K$), and $\mathcal{L}_{\text{swap}}$ enforces time-reversal symmetry at the midpoint for consistency and self-correction.
For each training sample, we draw a uniform timestep $t$ from $\text{Uniform}[t_{\text{eps}},1-t_{\text{eps}}]$, with $t_{\text{eps}}=1\times 10^{-3}$. We compute the forward prediction at $t$ for the main loss term, and evaluate it once more at the complementary timestep 1t1-t; this time-reversed prediction at 1t1-t is used to compute the swap loss.
4.3 Endpoint Anchoring and Boundary Behavior
[…]
Figure 6: Endpoint anchoring and boundary behavior diagnostics on AFHQ held-out dataset, averaged across classes.
4.4 Sampling
GAF supports three sampling modes: IER (Section 4.4.1), ODE (Section 4.4.2), and hybrid (Section 4.4.3).
4.4.1 Iterative Endpoint Refinement Sampling
Given the GAF factorization of its endpoint $J$ and $K$, we sample by iteratively refining endpoints on a linear bridge. In the forward direction, $K$ is used to refine the data endpoint from the intermediate bridge points $\mathbf{x_{t}}$. In the reverse direction, $J$ is used to refine the noise endpoint from intermediate bridge points. Both the bridge $\mathbf{x_{t}}=(1-t)\mathbf{z_{0}}+t\mathbf{x_{1}}$ and the endpoints update ($\mathbf{x}_{1}\leftarrow(1-\alpha)\,\mathbf{x}_{1}+\alpha\,K$ for data, $\hat{\mathbf{z}}\leftarrow(1-\alpha)\,\hat{\mathbf{z}}+\alpha\,J$ for noise) are linear operations, so IER traverses straight paths between noise and data without ODE integration.
Forward Generation (noise $\to$ data). Starting from $\mathbf{z}_{y}\sim\mathcal{N}(0,I)$, we initialize $\mathbf{x_{1}}\leftarrow\mathbf{z_{0}}$, set $\alpha\in[0,1]$, and iterate over $t~\in[t_{\varepsilon},1-t_{\varepsilon}]$ (Algorithm 1). At each step, we form $\mathbf{x_{t}}=(1-t)\mathbf{z_{0}}+t\mathbf{x_{1}}$, predict $K$, and update $\mathbf{x_{1}}$ by a convex combination of the current data estimate and the prediction.
Reverse Inversion (data $\to$ noise). Given the data anchor $\mathbf{x_{\text{anchor}}}$, we initialize $\hat{\mathbf{z}}\sim\mathcal{N}(0,I)$, set $\alpha\in[0,1]$, and iterate over $t~\in[t_{\varepsilon},1-t_{\varepsilon}]$ (Algorithm 2). At each step, we form $\mathbf{x_{t}}=(1-t)\hat{\mathbf{z}}+t\mathbf{x_{\text{anchor}}}$, predict $J$, and update $\hat{\mathbf{z}}$ by a convex combination of the current noise estimate and the prediction.
Input: noise $\mathbf{z}_{0}$, steps $S$, $t_{\varepsilon}$, $\alpha$, $\mathrm{GAF}(\cdot,t)$data anchor $\mathbf{x}_{\mathrm{anchor}}$
Output: $\mathbf{x}_{1}$ (data estimate)
$\mathbf{x}_{1}\leftarrow\mathbf{z}_{0}$
$t_{1},\dots,t_{S}\leftarrow\mathrm{Linspace}(t_{\varepsilon},\,1-t_{\varepsilon},\,S)$
for $s\leftarrow 1$ to $S$ do
 $t\leftarrow t_{s}$
 $\mathbf{x}_{t}\leftarrow(1-t)\,\mathbf{z}_{0}+t\,\mathbf{x}_{1}$
 $(\_,K)\leftarrow\mathrm{GAF}(\mathbf{x}_{t},t)$
 $\mathbf{x}_{1}\leftarrow(1-\alpha)\,\mathbf{x}_{1}+\alpha\,K$
end for
return $\mathbf{x}_{1}$
Algorithm 1 IER Forward (noise $\rightarrow$ data; $K$-only)
Input: data anchor $\mathbf{x}_{\mathrm{anchor}}$, steps $S$, $t_{\varepsilon}$, $\alpha$, $\mathrm{GAF}(\cdot,t)$
Output: $\hat{\mathbf{z}}$ (noise estimate)
$\hat{\mathbf{z}}\sim\mathcal{N}(0,I)$
$t_{1},\dots,t_{S}\leftarrow\mathrm{Linspace}(1-t_{\varepsilon},\,t_{\varepsilon},\,S)$
for $s\leftarrow 1$ to $S$ do
 $t\leftarrow t_{s}$
 $\mathbf{x}_{t}\leftarrow(1-t)\,\hat{\mathbf{z}}+t\,\mathbf{x}_{\mathrm{anchor}}$
 $(J,\_)\leftarrow\mathrm{GAF}(\mathbf{x}_{t},t)$
 $\hat{\mathbf{z}}\leftarrow(1-\alpha)\,\hat{\mathbf{z}}+\alpha\,J$
end for
return $\hat{\mathbf{z}}$
Algorithm 2 IER Reverse (data $\rightarrow$ noise; $J$-only)
4.4.2 ODE Sampling
At inference, we generate samples by integrating the emergent velocity field v=KJv=K-J from noise ($t=0$) to data ($t=1$) using Euler’s method [15]. Unlike standard models, GAF defines this field algebraically, enabling precise compositional generation directly within the ODE solver. As during training, we discretize the time interval $t_{\text{eps}},1-t_{\text{eps}}$] into uniformly spaced steps:
$t_{0}=t_{\text{eps}},t_{1},\dots,1-t_{\text{eps}},\text{with}~t_{\text{eps}}=1\times 10^{-3}$
At each step we update the latent state via
$z_{k+1}=z_{k}+(t_{k+1}-t_{k})v(z_{k},t_{k}).$ (32)
We initialize $\mathbf{z}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ at $t=t_{\text{eps}}$ and evolve to $t=1-t_{\text{eps}}$. For multi-class generation, samples are conditioned on their target class label $y$ through the corresponding $K_{n}$ heads.
4.4.3 Hybrid Sampling
GAF supports mid-trajectory switching between IER and ODE at any step, in both forward (generation) and reverse (inversion) directions. Run Algorithm 1 until $t_{s}$, then continue with Algorithm 2, or vice versa.
5 Experimental Evaluation
In this section, we evaluate GAF for image generation, as proposed in Sections 3 and 4. We first describe the dataset (Section 5.1.1) and training setup (Section 5.1.2), then present results for single-class (Section 5.2) and multi-class generation (Section 5.3), followed by qualitative evaluation (Section 5.4), an ablation study (Section 5.5), transport algebra (Section 5.6), and a comparison with Rectified Flow (Section 5.7).
5.1 Experimental Setup
5.1.1 Dataset
We train and evaluate GAF on four datasets.
CIFAR-10 [18] contains 50,000 training images across 10 object classes at $32\times 32$ px resolution, with $5,000$ images per class. We use this dataset for a $10$-class (i.e., one $J$ and ten $K$) generation task.
CelebA-HQ [25] consists of $202,599$ celebrity face images resized to $256\times 256$ px resolutions. We use this dataset for a single-class (i.e., one $J$ and one $K$) generation task.
AFHQ [5] contains $15,000$ high-quality images of cats, dogs, and wildlife at $512\times 512$ px resolution. We use this dataset for a $3$-class (i.e., one $J$ and three $K$) generation task.
ImageNet-1k [6, 33] is the full ImageNet classification dataset with $1.28$ million training images across $1,000$ object classes, with $1,300$ images per classes. For demonstration purposes, we use a single-class dataset (school bus). Additionally, we use a multi-class subset consisting of $1,000$ distinct classes. For this multi-class generation, we use $1,000$ independent $K$ heads, each corresponding to a specific class.
5.1.2 Training
Through an extensive parameter sweep of the residual and swap loss coefficients, we found $\lambda_{\text{res}}=0.003$ and $\lambda_{\text{swap}}=0.002$ to perform best. We use these values in all experiments unless stated otherwise. The training images are first encoded by the SD1.5-VAE [32]. We use the AdamW optimizer with $\beta=(0.9,0.99)$, and learning rate $1\times 10^{-4}$ for all of the datasets, with no weight decay in any of the training runs. The ImageNet-1k School Bus class is trained for $300,000$ iterations, whereas the other models are trained for $100,000$ iterations.
For the ImageNet multi-class experiment, we initialize GAF’s trunk from pretrained DiT-XL/2 weights and train the trunk and the $J/K$ heads jointly. We refer to this as retrunking: repurposing an existing pretrained backbone as GAF’s shared trunk. The trunk and endpoint heads are architecturally independent, so any compatible pretrained backbone can be adopted as the starting point for GAF training. The complete training configurations for each dataset is shown in Table 2.
Table 2: Training configurations for each dataset.
CIFAR-10 CelebA-HQ ImageNet
Training scratch scratch retrunked
Architecture DiT-B/2 DiT-XL/2 DiT-XL/2
VAE SD 1.5 SD 1.5
Batch size 256 32 128
Iterations 100k 100k 100k
Learning rate 2e-4 1e-4 1e-4
Warmup steps 500 3000 0
$\beta_{1},\beta_{2}$ 0.9, 0.99 0.9, 0.99 0.9, 0.99
EMA decay 0.9995 0.9999 0.9999
$\lambda_{\text{res}}$ 0.003 0.003 0.0001
$\lambda_{\text{time}}$ 0.002 0.002 0.0001
5.2 Single-Class Generation (One $J$ and One $K$)
This section validates GAF on single-class generation, using the school bus class from ImageNet, which contains about 1.3k images used for training. Figure 7 shows some example generations. We can see that a single K head model with 250-step Heun sampling demonstrates high-quality, diverse generations. Note that the generations showcase varied viewpoints, lighting conditions, and backgrounds while maintaining class consistency.
Figure 7: Single-class generated samples ($256\times 256$ px) from GAF trained on the ImageNet school bus class.
5.3 Multi-Class Generation (One $J$ and Multiple $K$s)
We evaluate GAF on multi-class generation on three datasets. First, we utilize CIFAR-$10$ ($32\times 32$ px) with ten object classes and one $J$ head and ten $K$ heads. Second, we utilize AFHQ ($512\times 512$ px) with three classes: cat, dog, and wild animals, using one $J$ head and three $K$ heads. Third, we utilize ImageNet-1k with $1,000$ classes, using one $J$ head and $1,000$ $K$ heads. Figure 8 and Figure 9 show samples from each independent $K$ head, each corresponding to one of the three or thousand classes with in each dataset, respectively. This demonstrates that GAF scales to multi-class generation while maintaining high per-class quality. Each $K$ head produces diverse, high-fidelity samples within its class without interference from other heads.
Figure 8: Multi-class generated samples ($512\times 512$ px) from GAF trained with one $J$ and three independent $K$ heads, on three classes (cat, dog, and wild) of the AFHQ dataset.
Figure 9: Comprehensive gallery of generated samples from the $1,000$ class ImageNet dataset. The figure demonstrates the model’s ability to handle high intra-class variance.
5.4 Qualitative Evaluation
We analyze the effect of the number of IER, Euler, and Heun steps on generation quality using CelebA-HQ ($256\times 256$) and ImageNet ($256\times 256$). We quantitatively measure sample quality with the Fréchet Inception Distance (FID) [12, 37] and inception score (IS) [35], and the diversity using the recall score [19]. All FID scores are computed by the PyTorch Fidelity library [29] using 50,000 generated samples.
Figure 10: Effect of Euler steps on CelebA-HQ $256$ sample quality.
Figure 11: Effect of IER steps on CelebA-HQ $256$ sample quality.
As shown in Figure 10, with very few Euler steps (151-5) the samples are blurry but roughly coherent; increasing the steps to 5405-40 significantly improves the quality and semantic detail. The FID (calculated using Heun, Figure 12a) value decreases rapidly up to about $40$ steps, reflecting improved sample coherence, and then flattens, with strong diminishing returns beyond roughly $80$ steps. Beyond $40$ steps, the curve exhibits clear diminishing returns: FID improves from about $7.45$ at $40$ steps to only $7.27$ at $80$ steps. Therefore, while $80$ steps yield the best score, high-quality generation is already achieved around 2020-$40$ Heun steps. The IER sampler, of which samples are presented in Figure 11, converges with far fewer steps: requiring only 686-8 steps for high-quality generation; and additional IER steps provide diminishing results.
Table 3 and 4 reports FID scores on ImageNet $256$, CIFAR-10, and CelebA-HQ $256$ using the Heun sampler. On ImageNet $256$, GAF achieves an FID of $7.51$ (at $N=250$ steps), outperforming DiT-XL/2 ($9.6$) and SiT-XL ($8.3$), with competitive IS, precision and recall. On CelebA-HQ $256$, GAF achieves an FID of $7.27$ (at $N=80$ steps), trailing LDM-4 ($5.11$) which employs heavily optimized diffusion pipeline. GAF achieves an FID of $9.53$ on CIFAR-10 dataset.
Table 3: FID results on ImageNet $256\times 256$. All results without classifier-free guidance.
ImageNet 256
[…]
DiT-XL/2 675 7M 9.6 121.50 0.67 0.67
SiT-XL/2 675 7M 8.3 131.65 0.68 0.67
GAF (ours, retrunked) 694 100k 7.51 114.81 0.54 0.69
Table 4: FID comparison on CelebeA-HQ $256\times 256$ and CIFAR-$10$
CelebA-HQ 256
[…]
Glow [17] 68.93
SDE [43] 7.23
LSGM [44] 7.22
LDM-4 [32] 5.11
GAF (ours) 7.27
CIFAR-10
[…]
DDPM 3.17
Score SDE 2.20
Flow Match. 6.35
GAF (ours) 9.53
Figure 12: Effect of the number of IER, and Heun steps on sample quality, measured by FID.
5.5 Ablation Study
The ablation study compares the full GAF objective, $\mathcal{L}_{\text{GAF}}=\mathcal{L}_{\text{pair}}+\lambda_{\text{res}}\,\mathcal{L}_{\text{res}}+\lambda_{\text{swap}}\,\mathcal{L}_{\text{swap}}$, against a pair-only variant trained with $\mathcal{L}_{\text{GAF}}=\mathcal{L}_{\text{pair}}$ alone (i.e, removing residual and swap losses).
5.5.1 Effect of Swap and Residual Loss
We report Pixel MSE, LPIPS(Alex), LPIPS(VGG), Latent MSE, and Latent LPIPS-GAF (computed from trunk features); lower is better for all metrics. The MSE captures the reconstruction error between the initial and final image, while LPIPS captures perceptual similarity. Latent space metrics measure transport consistency before VAE decoding.
Figure 13: Effect of swap and residual losses under cyclic transport ($z_{y}\rightarrow K_{0}\xrightarrow{J}K_{1}\xrightarrow{J}K_{2}\xrightarrow{J}K_{0}$).
As shown in Figure 13, across the cyclic transport, the full objective yields consistently low error rates and more stable multi-class transport at lower step counts. Removing $\mathcal{L}_{\text{res}}$ and $\lambda_{\text{swap}}$ does not collapse generation indicating that the $\mathcal{L}_{\text{pair}}$ loss is the dominant learning signal and the residual, and the time antisymmetric losses both serve as regulizers. At larger step counts ($N>100$), the error values of the full and pair objectives values converge. The $\mathcal{L}_{\text{res}}$ and $\lambda_{\text{swap}}$ are not required for endpoint learning; instead, they primarily act as stability and efficiency regularizers.
5.5.2 Image-to-Image Translation
Figure 14a shows an 8-step IER image-to-image translation using cross-modality transport between the cat and dog classes, as described in Section 3.3.2. The base images in Figure 14b are from the AFHQ held-out dataset. The images generated with GAF trained using the $\mathcal{L}_{\text{pair}}$-only objective (Figure 14c) show more drift and local artifact and inconsistent multi-class generation, while the full loss objective (Figure 14d) preserves identity and semantic structure more reliably under chained transport. This indicates that the residual and the swap losses are important for stabilization of the latent space.
Figure 14: Image-to-Image translation (8-step IER). (a) Visualization of the full cross-modality translation path ($K_{\text{cat}}\xrightarrow{J}K_{\text{dog}}$). (b) Base images from the held-out AFHQ dataset. (c) Translations generated by GAF trained only with pair loss ($\mathcal{L}_{\text{pair}}$), which exhibit significant drift and local artifacts. (d) Translations utilizing the full loss objective, demonstrating that the residual and swap losses are crucial for stabilizing the latent space and preserving semantic structure.
5.6 Transport Algebra
This section showcases three applications of GAF’s novel transport algebra: pairwise interpolation (in Section 5.6.1), continuous cyclic transport (in Section 5.6.2), and barycentric transport (in Section 5.6.4).
5.6.1 Pairwise Interpolation
We demonstrate direct $K$ head interpolation, by performing the operation $K_{\text{interp}}=(1-\alpha)K_{i}+\alpha K_{j}$ or $v=(1-\alpha)v_{i}+\alpha v_{j}$, for $\alpha\in[0,1].$ Figure 15 shows samples from interpolating between two classes from the AFHQ dataset, trained for multi-class generation in Section 5.3. We can see a smooth semantic transition between all three class pairs. Each interpolation produces coherent intermediate states, validating that independent $K$ heads enable linear composition directly over the K heads or in velocity space.
[…]
Figure 15: Pairwise interpolation samples ($512\times 512$ px) from GAF trained with one $J$ and three independent $K$ heads on the AFHQ dataset.
5.6.2 Continuous Cyclic Transport.
Cyclic Endpoint Transport. We quantitatively analyze GAF’s endpoint transport by running three RK4 [4, 15] integration steps on AFHQ dataset (Cat $\to$ Dog $\to$ Wild $\to$ Cat, Dog $\to$ Wild $\to$ Cat $\to$ Dog, and Wild $\to$ Cat $\to$ Dog $\to$ Wild) and measuring LPIPS [48], pixel MSE (post-VAE), and latent MSE (pre-VAE) between the initial and final states.
We use the cross-modality transport operator via $J$ defined in Section 3.3. Each cycle (e.g., Cat $\to$ Dog, Dog $\to$ Wild, Wild $\to$ Cat) is implemented as the encode–decode path: $v_{i}=K_{i}-J$, using the following method:
(A) Decode from the shared noise anchor $z_{y}$ to class $i$ with $v_{i}$ (forward in time),
(B) Encode back to $J$ with $v_{i}$ (reverse in time), and then,
(C) Decode to the next class $j$ with $v_{i\to j}$.
Figure 16: Cyclic Transport Analysis. We visualize the cyclic transport maps across classes (Top) and their corresponding convergence error rates (Bottom). The trajectories correspond to cyclic transport along the cycle $v_{i\to j\to k\to i}$ across 4 randomly chosen semantic cycles. The error difference between the first and the last image decreases monotonically, reaching machine precision ($\approx 10^{-16}$) at $N=1,000$.
Thus, a full Cat $\to$ Dog $\to$ Wild $\to$ Cat cycle is a sequence of decode–encode legs that always returns through $J$, as in the general cross-modality path $v_{i}\xrightarrow{-(v_{i})}J\xrightarrow{}v_{j}$. The metrics in Figure 16 report the difference between the first and last endpoint of each cycle, both in latent space (pre-VAE) and image space (post-VAE).
Cyclic Velocity Space Interpolation. We demonstrate that GAF enables continuous cyclic transport using cyclic velocity space interpolation, as shown in Figure 17. We use the same multi-class model as in Section 5.3. For a given pair of classes ($i,j$), we define a parametric family of interpolated data spaces or velocity fields:
$K_{i\to j}=(1-\alpha)K_{i}+\alpha K_{j},\quad\alpha\in[0,1],$
$v_{i\to j}=(1-\alpha)v_{i}+\alpha v_{j},\quad\alpha\in[0,1],$
where $v_{i}=K_{i}-J$ and $v_{j}=K_{j}-J$ are the class-specific velocity fields. The full cyclic interpolation between the three classes $i,j$, and $k$ is computed as:
$K_{i}\xrightarrow{J}K_{j}\xrightarrow{J}K_{k}\xrightarrow{J}K_{i}$
$v_{i}\xrightarrow{-v_{i}}J\xrightarrow{}v_{i\to j}\xrightarrow{-v_{i\to j}}J\xrightarrow{}v_{j\to k}\xrightarrow{-v_{j\to k}}J\xrightarrow{}v_{i}$
We sample $\alpha$ at $10$ uniformly spaced values in $[0,1]$ and, for each $\alpha$, refine the endpoint or integrate the velocity field by recovering the noise endpoint from the previous data endpoint using IER or the same ODE solver. Class $i$ is the source class and $j$ is the target class; the interpolated velocity field $v_{i\to j}$ moves from $i$-manifold toward the $j$-manifold while changing only the velocity field, not the underlying integration scheme or random seed. The velocity field $v_{i\to j}$ maintains semantic coherence throughout the transition across classes.
Across all three cycles and $3,000$ random $z_{0}$ samples, we observe an average latent LPIPS of $\approx$ $1.3\times 10^{-16}$ between the initial and the final images in each cycle ($N=1,000$). This indicates GAF’s transport algebra is deterministically reversible. The near-perfect closure validates that independent $K$ heads form a consistent algebraic structure where cyclic compositions return exactly to their starting point.
Figure 17 demonstrates continuous identity-preserving transport through multiple classes. Starting from an initial sample from the first class, we sequentially interpolate through two additional classes before returning to the original class, using velocity blending for each pairwise transition (as in Section 5.6.1). The endpoint of each interpolation becomes the starting point of the next, maintaining latent continuity throughout the cycle.
[…]
Figure 17: Continuous cyclic interpolation transport preserving identity across three classes. Each row shows a complete cycle starting from one class (left) to two interpolated classes, and finally returning to the initial class. Distinctive features (coloring, facial marking) are semantically preserved throughout each cycle while class-specific structure transforms.
5.6.3 Spatial Velocity Interpolation
Figure 1819, and 20 shows that spatial composition can be performed either in endpoint space using $K$ heads or in velocity space using $v$. Using region masks ($m$), we combine class-specific endpoint heads directly (e.g., ($K_{\text{cat}},K_{\text{dog}},K_{\text{wild}}$)) to assign different semantics to different facial regions, while keeping the remaining region fixed. For example, we can assign masks to target specific region by dividing the area into $N\in\mathbb{R}$ segments horizontally, diagonally, radially, or in any random configuration, as shown in Figure 20. We can also assign masks to regions such as ears (E), eyes (I), mouth (M), nose (N), and rest (R) (the remaining region not covered by the masks), as shown in Figure 18, to composite features that are directed by the IER or velocity space integration. This produces coherent localized edits that preserve global identity in non-targeted regions, indicating that endpoint composition is a native control mechanism in GAF. Endpoint mixing offers a direct coordinate-level generation (see Figure 19), while velocity mixing uses the ODE generation (see Figure 18 and 20).
Figure 18: Spatial composition on AFHQ using masked velocity blending.
Figure 19: Spatial composition on AFHQ using masked $K$ head blending (8-step IER).
Figure 20: Spatial composition on ImageNet using masked velocity blending.
5.6.4 Barycentric Transport
We extend pairwise interpolation to a three-way composition using barycentric coordinates. The blended velocity field is $v_{\text{blend}}=\alpha v_{i}+\beta v_{j}+\gamma v_{k}$, where $\alpha+\beta+\gamma=1$. We use square-to-simplex mapping with $\alpha=u(1-u),~\beta=(1-u)(1-v),$ and $\gamma=v$ to cover the full compositional space.
(a) Dog-cat-wild.
(b) Cat-dog-wild.
(c) Wild-dog-cat.
vv01uu01Class 1 ($\beta=1$)Class 0 ($\alpha=1$)Class 2($\gamma=1$)
(d) Interpolation grid map showing the mixing ratios.
Figure 21: Barycentric Transport Algebra. Visualization of three-way velocity blending across the Dog, Cat, and Wild classes. (a-c) $7\times 7$ grids showing the semantic simplex. Corners represent pure classes; interior points show linear interpolations of the vector fields. (d) The corresponding weight map.
Figure 21 shows three examples of three-way compositions, namely dog-cat-wild, cat-dog-wild, and wild-dog-cat. It shows all weighted combinations, with corners representing pure classes and the interior points showing multi-class blends.
5.7 Comparison with Rectified Flow
GAF learns endpoint operators $J$ and $K$ and samples via IER, while Rectified Flow learns a velocity field $v_{\theta}(\mathbf{x_{t}},t)$ and requires ODE integration for sampling. GAF has an equivalent velocity field v=KJv=K-J. Under ODE sampling, both models exhibit comparable compositional capacity.
Figure 22: GAF comparison with Rectified Flow under identical compositions. The first three images in the grid are the base images, while the fourth is the weighted blend.
Figures 22a, 22b, and 22c illustrate endpoint composition with IER and its qualitative comparison to ODE-based sampling. IER operates directly on the predicted coordinates without solving an ODE, a capability Rectified Flow’s velocity-only formalism cannot support. Because GAF exposes its two endpoints, $J$ and $K$, it also enables Transport Algebra operations directly in endpoint space and native bidirectional sampling, using $K$ for forward generation and $J$ for reverse inversion. Neither capability is present in velocity-only formalism. For example, one can compose endpoints as $0.5K_{\text{dog}}+0.3K_{\text{cat}}+0.2K_{\text{wild}}$ to generate a sample with mixed dog, cat, and wild features. Rectified Flow approximates this by constructing composite velocity field but must route through ODE integration. In contrast, GAF supports endpoint composition via IER, velocity composition via ODEs, and mid-trajectory switching between the two.
6 Conclusion
We introduced Generative Anchored Fields (GAF), a model that reframes generative modeling from direct trajectory prediction to endpoint coordinate learning. By learning endpoints $J$ and $K$ and deriving an emergent velocity field as v=KJv=K-J, GAF supports three inference methods within a single framework: endpoint iterative refinement (IER), ODE integration, and hybrid switching between them. Since GAF exposes its two endpoints, generation can be treated as algebra over learned transport coordinates. We demonstrated that GAF achieves FID $7.51$ on ImageNet $256\times 256$ without classifier-free guidance. GAF’s factorization into a trunk, $J$, and $K$ yeilds inherent modularity such that each component can be trained, frozen, or swapped independently. Since GAF separates representation learning (trunk) and transport parametrization (JJ/$K$ heads), pretrained vision backbones can serve as GAF’s shared trunk, inheriting Transport Algebra through $J/K$ head attachment without the need to train from scratch.
7 Future Work
GAF’s algebraic structure naturally extends to sequential domains. For video generation, we are reformulating Transport Algebra as Motion Algebra, where vector operations define transitions between frames. This will enable precise manipulation of temporal dynamics, offering a method to generate stable, controllable motion through the same arithmetic principles used for static data.
8 Acknowledgement
This work was funded in part by the Research Foundation Flanders (FWO) under Grant G0A2523N, IDLab (Ghent University-imec), Flanders Innovation and Entrepreneurship (VLAIO), and the European Union.